### v3

# YOLO Evolution: A Comprehensive Benchmark and Architectural Review of YOLOv12, YOLO11, and Their Previous Versions

Abstract
This study presents a comprehensive benchmark analysis of various YOLO (You Only Look Once) algorithms. It represents the first comprehensive experimental evaluation of YOLOv3 to the latest version, YOLOv12, on various object detection challenges. The challenges considered include varying object sizes, diverse aspect ratios, and small-sized objects of a single class, ensuring a comprehensive assessment across datasets with distinct challenges. To ensure a robust evaluation, we employ a comprehensive set of metrics, including Precision, Recall, Mean Average Precision (mAP), Processing Time, GFLOPs count, and Model Size. Our analysis highlights the distinctive strengths and limitations of each YOLO version. For example: YOLOv9 demonstrates substantial accuracy but struggles with detecting small objects and efficiency whereas YOLOv10 exhibits relatively lower accuracy due to architectural choices that affect its performance in overlapping object detection but excels in speed and efficiency. Additionally, the YOLO11 family consistently shows superior performance maintaining a remarkable balance of accuracy and efficiency. However, YOLOv12 delivered underwhelming results, with its complex architecture introducing computational overhead without significant performance gains. These results provide critical insights for both industry and academia, facilitating the selection of the most suitable YOLO algorithm for diverse applications and guiding future enhancements.
Index Terms:
YOLO (You Only Look Once), YOLOv12, YOLO11, Object detection, Ultralytics, Benchmark Analysis.
I Introduction
Object detection is an essential component of computer vision systems, enabling automated identification and localization of objects within images or video frames [37]. Its applications span from autonomous driving and robotics [17, 5, 21] [59] to inventory management, video surveillance, and sports analysis [4, 24, 58, 74].
Over the years, object detection has developed significantly. Initially, traditional methods such as the Viola-Jones algorithm [67] and the Deformable Part-based Model (DPM) [16] used handcrafted features and were effective for applications such as face detection [67], pedestrian detection [11], and video surveillance [3]. However, these methods had limitations in robustness and generalization. With the advancement of deep learning, network-based methods have since become the primary approach. These methods are usually classified into two categories: one-stage and two-stage approaches.
One-stage methods such as RetinaNet [34] and SSD (Single Shot MultiBox Detector) [38] perform detection in a single pass, balancing speed and accuracy. In contrast, two-stage methods, such as Region-based Convolutional Neural Networks (R-CNN) [20], generate region proposals and then perform classification, offering high precision but being computationally intensive.
Among one-stage object detection methods, YOLO (You Only Look Once) stands out for its robustness and efficiency. Initially introduced in 2015 by Redmon et al. [22], YOLO redefined object detection by predicting bounding boxes and class probabilities directly from full images in a single evaluation [50]. This innovative approach allowed YOLOv1 to achieve real-time object detection with impressive accuracy. Building upon this foundation, YOLOv2 [51] incorporated several key enhancements. It integrated the Darknet-19 framework, a 19-layer convolutional neural network that improved feature extraction. YOLOv2 also introduced batch normalization and employed data augmentation techniques inspired by the VGG architecture [60] to enhance the model’s generalization. YOLOv3 [52] further advanced the model with the Darknet-53 framework, a deeper network that significantly improved feature extraction capabilities. This version also utilized a Feature Pyramid Network (FPN)-inspired design, which allowed for better detection across various object scales by combining high-level semantic features with low-level detailed features, and a Three-Scale detection mechanism that improved accuracy for objects of different sizes.
Following YOLOv3, the model’s development branched into various communities, leading to several notable iterations. YOLOv4 [7], developed by Bochkovskiy et al., introduced enhancements such as Spatial Pyramid Pooling (SPP) and the Path Aggregation Network (PAN). SPP aggregates features from multiple scales, preserving spatial information, while PAN improves the fusion of features between layers, resulting in improved speed and accuracy. YOLOv5 [64] marked a significant transition by moving from the Darknet framework to PyTorch, a popular deep learning library. This transition made the model more accessible and easier to customize. The architecture incorporated strided convolution layers, which reduced computational load, and Spatial Pyramid Pooling Fast (SPPF) layers, optimizing memory usage while maintaining high performance. YOLOv6 and YOLOv7 continued this trajectory with innovative architectures. YOLOv6 [31] introduced RepVGG, an architecture that simplified convolutional layers during inference, and CSPStackRep blocks, which improve accuracy by splitting the feature map into two parts to process them separately. In addition, YOLOv6 employed a hybrid channel strategy for better feature representation. YOLOv7 [69] introduced the Extended Efficient Layer Aggregation Network (E-ELAN) to improve efficiency and effectiveness by enhancing information flow between layers.
Figure 1: Evolution of YOLO Algorithms throughout the years.
The most recent versions of YOLO, including YOLOv8, YOLOv9, YOLOv10, YOLO11, and YOLOv12 represent the forefront of the model’s development. YOLOv8 [61], released by Ultralytics, introduced semantic segmentation capabilities, allowing the model to classify each pixel of an image, and provided scalable versions to meet various application needs, from resource-constrained environments to high performance systems alongside other tasks such as pose estimation, image classification, and oriented object detection (OOB). YOLOv9 [70] built on its predecessors’ architectural advancements with Programmable Gradient Information (PGI), which optimizes gradient flow during training, and the Generalized Efficient Layer Aggregation Network (GELAN), which further improved performance by enhancing layer information flow. YOLOv10 [68], developed by Tsinghua University, eliminated the need for Non-Maximum Suppression (NMS) used by its predecessors, a technique used to eliminate duplicate predictions and pick the bounding boxes with the most confidence, by introducing a dual assignment strategy in its training protocol. Additionally, YOLOv10 features lightweight classification heads, spatial-channel decoupled downsampling, and rank-guided block design, making it one of the most efficient and effective YOLO models to date. YOLO11 [28], also introduced by Ultralytics, retains the capabilities of YOLOv8 with applications such as Instance Segmentation, Pose Estimation, and Oriented Object Detection while providing 5 scalable versions for different use cases. YOLO11 replaces the C2f block from YOLOv8 with the more efficient C3k2 block, delivering improved performance without compromising speed. Additionally, it introduces the C2PSA (Cross Stage Partial with Spatial Attention) module, which improves spatial attention in feature maps, increasing accuracy, especially for small and overlapping objects. Lastly, YOLOv12 joins the YOLO family by also implementing applications such as Pose Estimation, Instance Segmentation, and Oriented Object Detection. Additionally, it provides 5 scalable version and introduces new architectural advancements such as the Area Attention (A2) and the Residual Efficient Layer Aggregation Networks (R-ELAN), which improve training stability and model convergence through block-level residual design and optimized feature aggregation.
This object detection algorithm has undergone several developments, as seen in Figure reffig:yoloevolution achieving competitive results in terms of accuracy and speed, making it the preferred algorithm in various fields such as ADAS (Advanced Driver-Assist System) [50], video surveillance [41], face detection [42], and many more [19]. For instance, YOLO plays a crucial role in the agriculture field by being implemented in numerous applications such as crop classification [1] [18], pest detection [36], automated farming [71] [40], and virtual fencing [66]. Moreover, YOLO has been utilized on numerous occasions in the field of healthcare such as cancer detection [48], ulcer detection [2], medicine classification [39] [45], and health protocols enforcement [12].
In recent years, Ultralytics has played a crucial role in the advancement of YOLO by maintaining, improving, and making these models more accessible [49]. Notably, Ultralytics has streamlined the process of fine-tuning and customizing YOLO models, a task that was considerably more complex in earlier iterations. The introduction of user-friendly interfaces, comprehensive documentation, and pre-built modules has greatly simplified essential tasks such as data augmentation, model training, and evaluation. Moreover, the development of scalable model versions allows users to select models tailored to specific resource constraints and application requirements, thereby facilitating more effective fine-tuning. For instance, YOLOv8n is favorable over YOLOv8m in scenarios where speed and computational efficiency are prioritized over accuracy, making it ideal for resource-constrained environments. The integration of advanced tools for hyperparameter tuning, automated learning rate scheduling, and model pruning has further refined the customization process. Continuous updates and robust community support helped YOLO models to be more accessible and adaptable for a wide range of applications.
This paper aims to present a comprehensive comparative analysis of the YOLO algorithm’s evolution. It makes a significant contribution to the field by offering the first comprehensive evaluation of YOLOv12, the newest member of the YOLO family. By leveraging pre-trained models and fine-tuning them, we evaluate their performance across three diverse custom datasets, each with varying sizes and objectives. Consistent hyperparameters are applied to ensure a fair and unbiased comparison. The analysis focuses on numerous performance metrics, including speed, efficiency, accuracy, and computational complexity, as measured by GFLOPs count and model size. In addition, we explore the real-world applications of each YOLO version, highlighting their strengths and limitations across different use cases. Through this comparative study, we aim to provide valuable insights for researchers and practitioners, offering a deeper understanding of how these models can be effectively applied in various scenarios.
The rest of this paper is organized as follows: Section 2 covers related work. Section 3 describes the datasets, the models, and the experimental setup, including the hyperparameters and evaluation metrics used. Section 4 discusses the experimental results and comparative analysis. Finally, Section 5 concludes with insights drawn from the study.
II Related Work
The YOLO (You Only Look Once) algorithm is considered one of the most prominent object detection algorithms. It achieves state-of-the-art speed and accuracy, and its various applications have made it indispensable in numerous fields and industries. Numerous researchers have shown interest in this object detection algorithm by publishing papers reviewing its evolution, fine-tuning its models, and benchmarking its performance against other computer vision algorithms. This widespread interest underscores YOLO’s important role in advancing the field of computer vision.
The paper [15] examines seven semantic segmentation and detection algorithms, including YOLOv8, for cloud segmentation from remote sensing imagery. It conducts a benchmark analysis to evaluate their architectural approaches and identify the most performing ones based on accuracy, speed, and potential applications. The research aims to produce machine learning algorithms with cloud segmentation using a few spectral bands, including RGB and RGBN-IR combinations.
Showcasing the implementation of YOLO in marine ship detection, the study [25] propose the Adaptive Multi-scale YOLO (AM YOLO) to tackle marine ship segmentation challenges, introducing modules for enhanced detail extraction and cross-scale integration. While achieving up to 4.0% precision gains over YOLOv5s, the study focuses primarily on marine environments, limiting generalization to other domains—a gap addressed in this paper through broader dataset evaluations and versatile benchmarks alongside implementing newer state-of-the-art YOLO models.
In the paper [72], the authors developed a lightweight YOLO model optimized from YOLOv5, designed for multi-target garbage detection in dense and resource-constrained scenarios including MobileViTv3 for feature extraction, optimized MobileViT blocks for enhanced discrimination, and feature reuse techniques to minimize information loss. However, their reliance on YOLOv5 as a base architecture, which is less efficient and accurate than newer versions like YOLOv10 or YOLO11, limits both computational efficiency and performance—gaps addressed in this paper by leveraging state-of-the-art YOLO models and broader domain evaluations.
In the field of Smart Cities, the authors [6] work on a comprehensive survey of 3D object detection methods in road scene understanding. Their systematization offers a practical framework for comparing approaches, aiding researchers and practitioners in navigating the field. However, the study primarily focuses on other object detection models neglecting the use of YOLO in 3D object detection.
The paper [35] introduce the Efficient Attention Pyramid Transformer (EAPT), which enhances vision tasks by using Deformable Attention and the Encode-Decode Communication module (En-DeC) to capture more dynamic, global attention. While EAPT improves performance in various vision tasks, its integration with YOLO models is not explored. Incorporating these techniques into YOLO could enhance its feature extraction and global context understanding further improving the performance of YOLO variants. However, the study neglects the implementation of YOLO as a backbone, overlooking its potential in this area.
The authors of [23] review the evolution of the YOLO variants from version 1 to version 8, examining their internal architecture, key innovations, and benchmarked performance metrics. The paper highlights the models’ applications across domains like autonomous driving and healthcare, and proposes incorporating federated learning to improve privacy, adaptability, and generalization in collaborative training. The review, however, limits its focus to mAP (mean Average Precision) for accuracy evaluation, neglecting other key metrics such as Recall and Precision. Additionally, it considers FPS (frames per second) as the sole measure of computational efficiency, excluding the impact of preprocessing, inference, postprocessing times, GFLOPs, and size.
The paper [13] thoroughly analyzes single-stage object detectors, particularly YOLOs from YOLOv1 to YOLOv4, with updates to their architecture, performance metrics, and regression formulation. Additionally, it provides an overview of the comparison between two-stage and single-stage object detectors and applications utilizing two-stage detectors. However, not including newer YOLO models limits its comprehensiveness, leaving a gap in understanding the advancements and improvements introduced in more recent versions.
The authors of [56] explore the evolution of the YOLO algorithms from version 1 to 10, highlighting their impact on automotive safety, healthcare, industrial manufacturing, surveillance, and agriculture. The paper highlights incremental technological advances and challenges in each version, indicating a potential integration with multimodal, context-aware, and General Artificial Intelligence systems for future AI-driven applications. However, the paper does not include a benchmarking study or a comparative analysis of the YOLO models, leaving out performance comparisons across the versions.
The paper in [65] explores the development of the YOLO algorithm till the fourth version. It highlights its challenges and suggests new approaches, highlighting its impact on object detection and the need for ongoing study. However, it does not include newer YOLO models and lacks a comparative benchmark, limiting its ability to assess the progress and performance differences across various versions.
The authors in the work in [29] analyze the YOLO algorithm, focusing on its development and performance. They conduct a comparative analysis of the different versions of YOLO till the 8th version, highlighting the algorithm’s potential to provide insights into image and video recognition and addressing its issues and limitations. The paper focuses exclusively on the mAP metric, overlooking other accuracy measures such as Precision and Recall. Additionally, it neglects speed and efficiency metrics limiting the scope of the comparative study. The paper also omits the evaluation of the most recent models, YOLOv9, YOLOv10, and YOLO11.
This paper makes several key contributions: (i) It pioneers a comprehensive comparison of YOLOv12 against its predecessors across their scaled variants from nano- to extra-large; (ii) It provides deep insights into the architectural advancements of YOLO by analyzing key structural developments across versions. (iii) It evaluates YOLO models using three diverse datasets, reflecting various object properties and real-world applications, including Smart Cities, Satellite Imaging, and Wildlife Conservation. (iv) The performance evaluation extends beyond mAP and FPS, incorporating critical metrics such as Precision, Recall, Preprocessing Time, Inference Time, Postprocessing Time, GFLOPs, and model size. (v) These comprehensive metrics provide valuable insights for selecting the most optimal YOLO models, benefiting both industry professionals and academics. (vi) It offers specific use case recommendations, identifying the most suitable YOLO models for different scenarios and environments, such as resource-constrained deployments, real-time applications, and the detection of small or overlapping objects.
III Benchmark Setup
III-A Datasets
This study aims to conduct in-depth benchmark research and assess the YOLO algorithms provided by the Ultralytics library. The main goal is to provide a thorough and comparative analysis of these models and explain their strengths, deficiencies, and possible applications.
This paper is made possible using several publicly accessible datasets on Kaggle and Roboflow. The selection of the datasets is based on the increasing implementation of the YOLO algorithms in the fields of Autonomous driving [50] [57] [32] [8], satellite imagery [33] [9] [47], and wildlife conservation [53] [73] [54]. Moreover, each picked dataset presents unique difficulties and situations for object detection with varying image sizes and number of observations alongside the number of classes to ensure a comprehensive evaluation.
III-A1 Traffic Signs Dataset
The Traffic Signs dataset by Radu Oprea is an open-source dataset on Kaggle containing around 55 classes across 3253 training and 1128 validation images of traffic signs in various sizes and environments [43]. All images in the dataset are initially sized $\times×640$, with no labels for False Positives detection. To balance the classes, undersampling techniques were applied, as shown in Figure 2. After preprocessing the dataset by removing classes with low observation counts, 24 classes remained, with 3233 images split into 70% training, 20% validation, and 10% testing, with no data augmentation techniques applied. This dataset is vital for applications in autonomous driving, traffic management, road safety, and intelligent transportation systems. Additionally, it presents challenges due to the varying sizes of target objects and the similarities in patterns across different classes, which complicate the detection process.
Figure 2: Classes Distribution of the Traffic Signs Dataset
III-A2 Africa Wild Life Dataset
The Africa Wildlife dataset is an open-source Kaggle dataset by Bianca Ferreira, designed for real-time animal detection in nature reserves [62]. It features four common African animal classes: Buffalo, elephant, rhino, and zebra. Each class is represented by at least 376 images, which were collected via Google image searches without any data augmentation techniques applied. The dataset is split into 70% training, 20% validation, and 10% testing, with all images manually labeled in the YOLO format, as shown in Figure 3. The dataset presents challenges such as varying aspect ratios, with each image containing at least one instance of the specified animal class and potentially multiple occurrences of other classes. Furthermore, the overlapping of target objects increases the difficulty of detection. This dataset is crucial for applications in wildlife conservation, anti-poaching efforts, biodiversity monitoring, and ecological research.
Figure 3: Classes Distribution of the Africa Wildlife Dataset
III-A3 Ships/Vessels Dataset
The Ships/Vessels dataset is an extensive open-source collection containing approximately 13.5k images, collected by Siddharth Sah from numerous Roboflow datasets, curated explicitly for ship detection [55]. Each image has been manually annotated with bounding boxes in the YOLO format, ensuring precise and efficient detection of ships. This dataset features a single class, ”ship,” and is divided into 70% training, 20% validation, and 10% testing allowing for streamlined and focused analysis. Notably, no data augmentation techniques were applied during the experiment, which ensures that the model’s performance is evaluated on the raw, unaltered dataset. However, the relatively small size of the target objects and their varying rotations pose challenges for detection, particularly for the YOLO algorithm, which often struggles with small object detection and objects with varying orientations. The dataset is essential for various practical applications such as maritime safety, fisheries management, marine pollution monitoring, defense, maritime security, and more.
III-B Models
III-B1 Ultralytics vs. Original YOLO
In this subsection, we will conduct a comparative analysis between the models provided by Ultralytics and their original counterparts on the Traffic Signs dataset provided by Radu Oprea [43] using the same hyperparameters in Table V. The objective is to highlight the differences between Ultralytics models and the original versions, which justifies the exclusion of YOLOv4 [7], YOLOv6 [31], and YOLOv7 [69] from this paper due to the lack of support for these models by Ultralytics. This analysis will demonstrate why focusing exclusively on Ultralytics-supported models ensures a fair and consistent benchmark evaluation.
Ultralytics Supported Models and Tasks:
Ultralytics library provides researchers and programmers with various YOLO models for inference, validation, training, and export. Based on the results of Table I, we notice that Ultralytics does not support YOLOv1, YOLOv2, YOLOv4, and YOLOv7. Concerning YOLOv6, the library only supports the configuration *.yaml files without the pre-trained *.pt models.
TABLE I: Ultralytics-supported library tasks and models
YOLO
Version
Inference Validation Training
YOLOv1 No No No
YOLOv2 No No No
YOLOv3u Yes Yes Yes
YOLOv4 No No No
YOLOv5u Yes Yes Yes
YOLOv6 Yes Yes Yes
YOLOv7 No No No
YOLOv8 Yes Yes Yes
YOLOv9 Yes Yes Yes
YOLOv10 Yes Yes Yes
YOLO11 Yes Yes Yes
YOLOv12 Yes Yes Yes
00footnotetext: Boldface text shows available pre-trained models
Performance Comparison of Ultralytics and Original Models:
Based on the results of our comparative analysis of the Ultralytics models and their original counterparts on the Traffic Signs dataset presented in Table II, we observe significant discrepancies between the performance of the Ultralytics models and their original counterparts. Notably, Ultralytics’ versions of YOLOv5n (nano) and YOLOv3 demonstrate superior performance, underscoring the enhancements and optimizations implemented by Ultralytics. Conversely, the original YOLOv9c (compact) slightly outperforms its Ultralytics version, potentially due to the lack of extensive optimization for this newer model by Ultralytics. These observations highlight that the Ultralytics models have undergone substantial modifications, making a direct comparison with the original versions inequitable. Consequently, the noticeable performance discrepancy between the two models, including the original and Ultralytics models in the same benchmarking study, would not provide a fair or accurate assessment. Therefore, this paper will focus exclusively on the Ultralytics-supported versions to ensure consistent and fair benchmarks.
In total, 33 models from 7 different YOLO versions were trained on 3 different datasets as seen in table III. Each YOLO version includes several scaled models, (e.g., YOLOv5u, YOLOv5un, YOLOv5us, and YOLOv5ux), with the suffixes denoting model size and complexity, such as ”n” for nano, ”s” for small, ”m” for medium, ”l” for large, ”x” for extra-large, ”t” for tiny, ”c” for compact, ”b” for balanced, and ”e” for extended. These models offer a variety of trade-offs between accuracy and inference speed, as discussed in the following sections.
TABLE II: Ultralytics and original YOLO performance comparison
Version Source mAP50 mAP50-95
YOLOv9c Ultralytics 0.845 0.748
Github 0.881 0.786
YOLOv5n Ultralytics 0.756 0.663
Github 0.429 0.367
YOLOv3 Ultralytics 0.766 0.67
Github 0.562 0.471
TABLE III: YOLO versions and scaled versions
Version Scaled Version
YOLOv3u YOLOv3u-tiny
YOLOv3u
YOLOv5u YOLOv5un (nano)
YOLOv5us (small)
YOLOv5um (medium)
YOLOv5ul (large)
YOLOv5ux (extra-large)
YOLOv8 YOLOv8n (nano)
YOLOv8s (small)
YOLOv8m (medium)
YOLOv8l (large)
YOLOv8x (extra-large)
YOLOv9 YOLOv9t (tiny)
YOLOv9s (small)
YOLOv9m (medium)
YOLOv9c (compact)
YOLOv9e (extended)
YOLOv10 YOLOv10n (nano)
YOLOv10s (small)
YOLOv10m (medium)
YOLOv10b (balanced)
YOLOv10l (large)
YOLOv10x (extra-large)
YOLO11 YOLO11n (nano)
YOLO11s (small)
YOLO11m (medium)
YOLO11l (large)
YOLO11x (extra-large)
YOLOv12 YOLO11n (nano)
YOLOv12s (small)
YOLOv12m (medium)
YOLOv12l (large)
YOLOv12x (extra-large)
III-B2 YOLOv3u
YOLOv3, based on its predecessors, aims to improve localization errors and detection efficiency, particularly for smaller objects. It uses the Darknet-53 framework, which has 53 convolutional layers and achieves double the speed of ResNet-152 [52]. YOLOv3 also incorporates elements from the Feature Pyramid Network (FPN), such as residual blocks, skip connections, and up-sampling, to enhance its ability to detect objects efficiently across varying scales, as seen in Figure 4. The algorithm generates feature maps at three distinct scales, down-sampling the input at factors of 32, 16, and 8, and uses a three-scale detection mechanism to detect large, medium, and small-sized objects using distinct feature maps. Despite its improvements, YOLOv3 still faces challenges in achieving precise results for medium and large-sized objects, so Ultralytics released YOLOv3u. YOLOv3u is an improved version of YOLOv3 that utilizes an anchor-free detection method used later in YOLOv8 and improves upon the accuracy and speed of YOLOv3, especially for medium and large-sized objects.
Figure 4: YOLOv3 architecture showcasing the residual blocks and the upsampling layers [10].
III-B3 YOLOv5u
YOLOv5, proposed by Glenn Jocher, transitions from the Darknet framework to PyTorch, retaining many improvements from YOLOv4 [64] [26] and utilizing CSPDarknet as its backbone. CSPDarknet is a modified version of the original Darknet architecture that incorporates Cross-Stage Partial connections by splitting feature maps into separate paths, allowing for more efficient feature extraction and reduced computational costs. YOLOv5 features a strided convolution layer with a large window size, aiming to reduce memory and computational costs, as showcased in Figure 5. Moreover, this version adopts the Spatial Pyramid Pooling Fast (SPPF) module to provide a multiscale representation of the input feature maps. The SPPF module works by pooling features at different scales and concatenating them, allowing the network to capture fine and coarse information. This helps recognize objects of various sizes more effectively. In addition, YOLOv5 implements several augmentations, such as Mosaic, copy-paste, random affine, MixUp, HSV augmentation, and random horizontal flip. YOLOv5 is available in five variants, varying in width and depth of convolution modules. Ultralytics is actively improving this model through YOLOv5u, adopting an anchor-free detection method and achieving better overall performance, especially on complex objects of different sizes.
Figure 5: Detailed architecture of YOLOv5 including the CSPDarknet Backbone, PANet Neck, and YOLO Layer Head [14].
III-B4 YOLOv8
Ultralytics has introduced YOLOv8, a significant evolution in the YOLO series, with five scaled versions [61] [27]. Alongside object detection, YOLOv8 also provides various applications such as image classification, pose estimation, instance segmentation, and oriented object detection (OOB). Key features include a backbone similar to YOLOv5, with adjustments in the CSPLayer, now known as the C2f module, which combines high-level features with contextual information for enhanced detection accuracy highlighted in Figure 6. YOLOv8 also introduces a semantic segmentation model called YOLOv8-Seg, which combines a CSPDarknet53 feature extractor with a C2F module, achieving state-of-the-art results in object detection and semantic segmentation benchmarks while maintaining high efficiency.
Figure 6: Detailed architecture of YOLOv8 showcasing the backbone’s multiple convolutional layers to extract hierarchical features, the Feature Pyramid Network (FPN) enhances detection at different scales. The network head performs final predictions, incorporating convolutional blocks and upsample blocks to refine features [30].
III-B5 YOLOv9
YOLOv9, developed by Chien-Yao Wang, I-Hau Yeh, and Hong-Yuan Mark Liao, uses the Information Bottleneck Principle and Reversible Functions to preserve essential data across the network’s depth, ensuring reliable gradient generation and improved model convergence and performance [70]. Reversible functions, which can be inverted without loss of information, are another cornerstone of YOLOv9’s architecture. This property allows the network to retain a complete information flow, enabling more accurate updates to the model’s parameters. Moreover, YOLOv9 offers five scaled versions for different uses, focusing on lightweight models, which are often under-parameterized and prone to losing significant information during the feedforward process.
Programmable Gradient Information (PGI) is a significant advancement introduced in YOLOv9. PGI is a method that dynamically adjusts the gradient information during training to optimize learning efficiency. By selectively focusing on the most informative gradients, PGI helps preserve crucial information that might be lost in lightweight models. This advancement ensures the model retains the essential features for accurate object detection, improving overall performance.
In addition, YOLOv9 incorporates Gradient Enhanced Lightweight Architecture Network (GELAN), a new architectural advancement designed to improve parameter utilization and computational efficiency as illustrated in Figure 7. GELAN achieves this by optimizing the computational pathways within the network, allowing for better resource management and adaptability to various applications without compromising speed or accuracy.
Figure 7: YOLOv9 architecture featuring CSPNet, ELAN, and GELAN modules. CSPNet optimizes gradient flow and reduces computational complexity via feature map partitioning. ELAN enhances learning efficiency by linearly aggregating features, and GELAN extends this concept by integrating features from various depths and pathways [70].
III-B6 YOLOv10
YOLOv10, developed by Tsinghua University researchers, builds upon previous models’ strengths with key innovations [68]. The architecture has an enhanced Cross Stage Partial Network (CSPNet) backbone for improved gradient flow and reduced computational redundancy, as seen in Figure 8. The network is structured into three main components: the backbone, the neck, and the detection head. The neck includes Path Aggregation Network (PAN) layers for effective multiscale feature fusion. PAN is designed to enhance information flow by aggregating features from different layers, enabling the network to better capture and combine details at various scales, which is crucial for detecting objects of different sizes. At the same time, the One-to-Many Head generates multiple predictions per object during training to provide rich supervisory signals and improve learning accuracy. Five scaled versions are available, from nano to extra-large.
For inference, the One-to-One Head generates a single best prediction per object, eliminating the need for Non-Maximum Suppression (NMS). By removing the need for NMS, YOLOv10 reduces latency and improves the post-processing speed. In addition, YOLOv10 includes NMS-Free Training, which uses consistent dual assignments to reduce inference latency, and a model design that optimizes various components from both efficiency and accuracy perspectives. This includes lightweight classification heads, spatial-channel decoupled downsampling, and rank-guided block design. In addition, the model incorporates large-kernel convolutions and partial self-attention modules to enhance performance without significant computational costs.
Figure 8: YOLOv10 architecture showcasing the dual label assignment strategy for improving accuracy. The PAN layer enhances feature representation alongside one-to-many head for regression and classification tasks and one-to-one head for precise localization [68].
III-B7 YOLO11
YOLO11 [28] is one of the latest additions to the YOLO series developed by Ultralytics, building upon the developments of its predecessors, especially YOLOv8. This iteration offers five scaled models from nano to extra large, catering to various applications. Like YOLOv8, YOLO11 includes numerous applications such as object detection, instance segmentation, image classification, pose estimation, and oriented object detection (OBB).
Key improvements in YOLO11 include the introduction of the Cross-Stage Partial with Self-Attention (C2PSA) module, as seen in Figure 9, which combines the benefits of cross-stage partial networks with self-attention mechanisms. This enables the model to capture contextual information more effectively across multiple layers, improving object detection accuracy, especially for small and colluded objects. Additionally, in YOLO11, the C2f block has been replaced by C3k2, a custom implementation of the CSP Bottleneck that uses two convolutions, unlike YOLOv8’s use of one large convolution. This block uses a smaller kernel, retaining accuracy while improving efficiency and speed.
Figure 9: YOLO11 architecture showcasing the new C3k2 blocks and the C2PSA module. [28] [46].
III-B8 YOLOv12
YOLOv12 [63] is the latest evolution in the YOLO series, introducing an attention-centric design that significantly enhances both speed and accuracy. It continues the trend of offering five scalable models (Nano to Extra Large), making it adaptable for a wide range of applications such as object detection, instance segmentation, and oriented object detection (OBB). As seen in Figure 10, YOLOv12 introduces the Area Attention (A2) module, which maintains a large receptive field while drastically reducing computational complexity, allowing the model to enhance speed without compromising accuracy. Additionally, it features Residual Efficient Layer Aggregation Networks (R-ELAN), which improve training stability and model convergence through block-level residual design and optimized feature aggregation.
Moreover, YOLOv12 uses FlashAttention to minimize memory access overhead, closing the speed gap with CNNs. It adjusts the MLP ratio from 4 to 1.2, enhancing runtime efficiency, and removes positional encodings for a cleaner, faster model without losing detection accuracy. YOLOv12 also reduces complexity by using a single R-ELAN block in the last stage of the backbone instead of stacking three attention/CNN blocks. It replaces linear layers with convolutional layers and batch normalization, maximizing computational efficiency and achieving state-of-the-art latency-accuracy trade-offs.
Figure 10: YOLOv12 architecture showcasing the new Area Attention (A2) module and Residual Efficient Layer Aggregation Networks (R-ELAN). [63]
III-C Hardware and Software Setup
Table IV presents the libraries and software versions used throughout this study. The experiments were conducted with Python 3.12, Ubuntu 22.04, CUDA 12.5, and cuDNN 8.9.7 for GPU acceleration. Ultralytics 8.2.55 was used for model training, while WandB 0.17.4 was utilized for tracking experiments. To ensure a fair comparison, similar hyperparameters were used across all models, as outlined in Table V. The key hyperparameters include 100 epochs, the AdamW optimizer, a batch size of 16, and an image size of (640, 640). The learning rate was initially set to 0.0001 and increased to 0.01, while the dropout rate was set to 0.15. The dataset was split into 70% for training, 20% for validation, and 10% for testing with no data augmentation techniques applied to any dataset. The experiments were carried out using 2 NVIDIA RTX 4090 GPUs, each with 16,384 CUDA cores.
TABLE IV: Software setup
Name Version
Python 3.12
Ubuntu 22.04
CUDA 12.5
cuDNN 8.9.7
Ultralytics 8.2.55
WandB 0.17.4
TABLE V: Table of parameters
Parameter Value
Epochs 100
Optimizer AdamW
Batch Size 16
Image Size (640, 640)
Initial & Final Learning rate (0.0001, 0.01)
Dropout rate 0.15
Augmentation Techniques None
Data Split (70, 20, 10)
III-D Metrics
This study evaluates the performance of YOLO models using three primary metrics: accuracy, computational efficiency, and size. The accuracy metrics include Precision, Recall, Mean Average Precision at an Intersection over Union (IoU) threshold of 0.50 (mAP50), and Mean Average Precision at an Intersection over Union (IoU) threshold of 0.50-0.95 (mAP50-95). Precision [44] measures the ratio of correctly predicted observations to the total predicted observations, thus highlighting the occurrence of False Positives. Conversely, Recall [44] measures the ratio of correctly predicted observations to all actual observations, thus emphasizing the occurrence of False Negatives. Both mAP50 and mAP50-95 [75] provide a comprehensive summary of Precision and Recall. While mAP50 calculates the Mean Average Precision at an IoU threshold of 0.50, mAP50-95 extends this calculation across multiple IoU thresholds from 0.50 to 0.95, with 0.05 step size.
Regarding computational efficiency metrics, Preprocessing Time, Inference Time, and Postprocessing Time will be utilized to evaluate the model’s speed. Preprocessing Time refers to the duration taken to prepare raw data for input into the model. Inference Time is the duration required for the model to process the input data and generate predictions. Postprocessing Time denotes the time needed to convert the model’s raw predictions into a final, usable format. These metrics were measured using a sample of images for testing after training the models. Additionally, the Giga Floating-Point Operations Per Second (GFLOPs) measure the computational power for the model training, reflecting its efficiency. In contrast, the size metric reflects the actual disk size of the model and the number of its parameters.
These metrics are essential for providing a comprehensive overview of YOLO models’ performance, allowing for effective comparison and evaluation. These metrics allows a thorough assessment of the accuracy and efficiency of different YOLO models, ensuring a robust benchmark for their performance and applications in various real-world scenarios.
IV Benchmark Results and Discussion
IV-A Results
IV-A1 Traffic Signs Dataset
TABLE VI: Evaluation results for the Traffic Signs Dataset.
Versions Precision Recall mAP50 mAP50-95
Preprocess
Time
Inference
Time
Postprocess
Time
Total
Time
GFLOPs Size
YOLOv3u 0.75 0.849 0.874 0.781 0.7 8.5 0.4 9.6 207.86 282.4
YOLOV3u tiny 0.845 0.667 0.772 0.682 1.4 0.7 0.3 2.4 24.44 19
YOLOv5un 0.805 0.679 0.749 0.665 0.6 6.6 0.4 7.6 5.65 7.1
YOLOv5us 0.85 0.777 0.827 0.744 0.5 7.8 0.4 8.7 18.58 23.9
YOLOv5um 0.849 0.701 0.83 0.744 1.1 9.5 0.4 11 50.54 64.1
YOLOv5ul 0.831 0.836 0.886 0.799 0.6 9.7 0.4 10.7 106.85 134.9
YOLOv5ux 0.863 0.795 0.867 0.777 1.1 9.8 0.4 11.3 195.2 246.3
YOLOv8n 0.749 0.688 0.777 0.689 0.6 6.8 0.4 7.8 6.55 8.1
YOLOv8s 0.766 0.788 0.806 0.718 0.6 7.8 0.4 8.8 22.59 28.6
YOLOv8m 0.838 0.805 0.845 0.763 1.6 9.1 0.4 11.1 52.12 78.9
YOLOv8l 0.771 0.789 0.853 0.767 0.6 9.2 0.4 10.2 87.77 165
YOLOv8x 0.902 0.744 0.874 0.78 0.6 9.4 0.4 10.4 136.9 257.7
YOLOv9t 0.792 0.748 0.812 0.731 0.5 10 0.4 10.9 4.93 7.7
YOLOv9s 0.763 0.81 0.828 0.75 0.6 11.1 0.4 12.1 15.33 26.8
YOLOv9m 0.864 0.796 0.864 0.784 1 12.1 0.4 13.5 40.98 76.7
YOLOv9c 0.827 0.807 0.852 0.769 1.3 11.6 0.4 13.3 51.8 102.6
YOLOv9e 0.819 0.824 0.854 0.764 0.8 16.1 0.4 17.3 117.5 189.4
YOLOv10n 0.722 0.602 0.722 0.64 1 0.8 0.2 2 5.59 8.3
YOLOv10s 0.823 0.742 0.834 0.744 1.2 1.1 0.2 2.5 15.9 24.7
YOLOv10m 0.834 0.843 0.88 0.781 1.2 2.4 0.2 3.8 32.1 63.8
YOLOv10b 0.836 0.764 0.859 0.765 1 3.1 0.2 4.3 39.7 98.4
YOLOv10l 0.873 0.807 0.866 0.771 1.1 3.8 0.2 5.1 50 126.8
YOLOv10x 0.773 0.854 0.88 0.787 1 6.3 0.2 7.5 61.4 170.4
YOLO11n 0.768 0.695 0.757 0.668 1.2 0.6 0.4 2.2 5.35 6.4
YOLO11s 0.819 0.758 0.838 0.742 1.2 1 0.4 2.6 18.4 21.4
YOLO11m 0.898 0.826 0.893 0.795 1.2 2.4 0.4 4 38.8 67.9
YOLO11l 0.862 0.839 0.889 0.794 1.2 3 0.4 4.6 49 86.8
YOLO11x 0.819 0.816 0.885 0.784 0.9 6.1 0.4 7.4 109 194.8
YOLOv12n 0.883 0.684 0.802 0.708 1.2 4.6 0.4 6.2 5.34 6.4
YOLOv12s 0.761 0.839 0.877 0.78 1.6 4.3 0.5 6.4 18.1 21.4
YOLOv12m 0.868 0.811 0.89 0.79 0.9 4.9 0.4 6.2 39 67.3
YOLOv12l 0.826 0.836 0.877 0.781 0.9 5.5 0.4 6.8 51.2 89.6
YOLOv12x 0.832 0.849 0.889 0.793 1 9.5 0.5 11 114 198.9
Table VI presents a comparative analysis of the YOLO algorithms’ performance on the Traffic Signs dataset, evaluated based on accuracy, computational efficiency, and model size. The Traffic Signs dataset is a medium-sized dataset with varied object sizes, making it favorable for benchmarking. The results highlight the effectiveness of YOLO models in detecting traffic signs, demonstrating a range of precision. The highest mAP50-95 was 0.799, while the lowest recorded precision was 0.64. On the other hand, the highest mAP50 is 0.893 while the lowest is 0.722. The substantial gap between the mAP50 and mAP50-95 results suggests that the models have difficulties in uniformly handling traffic signs with different sizes at higher thresholds, reflecting areas for potential improvement in their detection algorithms.
Accuracy:
As illustrated in Figure 11, YOLOv5ul demonstrates the highest accuracy, achieving a mAP50 of 0.866 and a mAP50-95 of 0.799. This is followed by YOLO11m with a mAP50-95 of 0.795 and YOLO11l with a mAP50-95 of 0.794. In contrast, YOLOv10n exhibits the lowest precision, with a mAP50 of 0.722 and a mAP50-95 of 0.64, closely followed by YOLOv5un with a mAP50-95 of 0.665, as evidenced by the data points in Figure 11.
Figure 11: mAP50 and mAP50-95 YOLO results on traffic signs dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
Figure 12: Precision vs. Recall based on size on traffic signs dataset, with larger circles indicating larger model sizes.
Precision and Recall:
Figure 12 elucidates the trade-off between precision and recall taking the size of the models into consideration. Models such as YOLO11m, YOLO10l, YOLOv12m, YOLOv9m, YOLOv5ux, and YOLO11l exhibit high precision and recall, specifically with YOLO11m achieving a precision of 0.898 and a recall of 0.826 while having a size of 67.9Mb, and YOLOv12m achieving a precision of 0.868 and a recall of 0.811 with a significantly smaller size (39 Mb). In contrast, smaller models such as YOLOv10n (precision 0.722, recall 0.602), YOLOv8n (precision 0.749, recall 0.688), and YOLO11n (precision 0.768, recall 0.695) underperform in both metrics. This underscores the superior performance of larger models on the Traffic Signs dataset. Moreover, the high precision (0.883) and low recall (0.684) of YOLOv12n indicate a propensity for false negatives, while YOLOv3u’s high recall (0.849) and low precision (0.75) suggest a tendency for false positives.
Figure 13: Total processing time and GFLOPs count results on traffic signs dataset.
Computational Efficiency:
In terms of computational efficiency, YOLOv10n is the most efficient, with a processing time of 2ms per image and a GFLOPs count of 8.3, as shown in Figure 13. YOLO11n closely trails this at 2.2ms with a 6.4 GFLOPs count, and YOLOv3u-tiny with a processing time of 2.4ms and a GFLOPs count of 19, making it relatively computationally inefficient compared to the other fast models. However, the data indicates that YOLOv9e, YOLOv9m, YOLOv9c, and YOLOv9s are the least efficient, with inference times of 16.1ms, 12.1ms, 11.6ms, and 11.1ms, and GFLOPs count of 189.4, 76.7, 102.6, and 26.8, respectively. These findings delineate a clear trade-off between accuracy and computational efficiency.
Overall Performance:
When evaluating overall performance, which includes accuracy, size, and model efficiency, YOLO11m emerges as a consistently top-performing model. It achieves a mAP50-95 of 0.795, an inference time of 2.4ms, a model size of 38.8Mb, and a 67.9 GFLOPs count, as detailed in Figures 11, 13, and Table VI. This is followed by YOLO11l (mAP50-95 of 0.794, inference time of 4.6ms, size of 49Mb, and 86.8 GFLOPs count), and YOLOv10m (mAP50-95 of 0.781, inference time of 2.4ms, size of 32.1Mb, 63.8 GFLOPs count). These results highlight the robustness of these models in detecting traffic signs of various sizes while maintaining short inference times and small model sizes.
On the contrary, YOLOv9e demonstrates poor performance overall, achieving a relatively low mAP50-95 of 0.764, an inference time of 16.1ms, a large model size of 117.5MB, and one of the highest GFLOPs counts at 198.4. Additionally, YOLOv5um, YOLOv8m, and YOLOv8s performed suboptimally, further indicating the inconsistency in performance among earlier model families.
Notably, the YOLO11 and YOLOv12 families significantly outperform other YOLO families in terms of accuracy and computational efficiency. Their models consistently surpass counterparts from the YOLOv3u, YOLOv5u, YOLOv8, YOLOv9, YOLOv10 families, demonstrating their ability to balance precision and speed effectively in the detection of objects with varying sizes.
IV-A2 Africa Wildlife Dataset
TABLE VII: Evaluation results for the Africa Wildlife dataset.
Versions Precision Recall mAP50 mAP50-95
Preprocess
Time
Inference
Time
Postprocess
Time
Total
Time
GFLOPs Size
YOLOv3u 0.75 0.849 0.874 0.781 0.7 8.5 0.4 9.6 207.86 282.4
YOLOV3u tiny 0.845 0.667 0.772 0.682 1.4 0.7 0.3 2.4 24.44 19
YOLOv5un 0.805 0.679 0.749 0.665 0.6 6.6 0.4 7.6 5.65 7.1
YOLOv5us 0.85 0.777 0.827 0.744 0.5 7.8 0.4 8.7 18.58 23.9
YOLOv5um 0.849 0.701 0.83 0.744 1.1 9.5 0.4 11 50.54 64.1
YOLOv5ul 0.831 0.836 0.886 0.799 0.6 9.7 0.4 10.7 106.85 134.9
YOLOv5ux 0.863 0.795 0.867 0.777 1.1 9.8 0.4 11.3 195.2 246.3
YOLOv8n 0.749 0.688 0.777 0.689 0.6 6.8 0.4 7.8 6.55 8.1
YOLOv8s 0.766 0.788 0.806 0.718 0.6 7.8 0.4 8.8 22.59 28.6
YOLOv8m 0.838 0.805 0.845 0.763 1.6 9.1 0.4 11.1 52.12 78.9
YOLOv8l 0.771 0.789 0.853 0.767 0.6 9.2 0.4 10.2 87.77 165
YOLOv8x 0.902 0.744 0.874 0.78 0.6 9.4 0.4 10.4 136.9 257.7
YOLOv9t 0.792 0.748 0.812 0.731 0.5 10 0.4 10.9 4.93 7.7
YOLOv9s 0.763 0.81 0.828 0.75 0.6 11.1 0.4 12.1 15.33 26.8
YOLOv9m 0.864 0.796 0.864 0.784 1 12.1 0.4 13.5 40.98 76.7
YOLOv9c 0.827 0.807 0.852 0.769 1.3 11.6 0.4 13.3 51.8 102.6
YOLOv9e 0.819 0.824 0.854 0.764 0.8 16.1 0.4 17.3 117.5 189.4
YOLOv10n 0.722 0.602 0.722 0.64 1 0.8 0.2 2 5.59 8.3
YOLOv10s 0.823 0.742 0.834 0.744 1.2 1.1 0.2 2.5 15.9 24.7
YOLOv10m 0.834 0.843 0.88 0.781 1.2 2.4 0.2 3.8 32.1 63.8
YOLOv10b 0.836 0.764 0.859 0.765 1 3.1 0.2 4.3 39.7 98.4
YOLOv10l 0.873 0.807 0.866 0.771 1.1 3.8 0.2 5.1 50 126.8
YOLOv10x 0.773 0.854 0.88 0.787 1 6.3 0.2 7.5 61.4 170.4
YOLO11n 0.768 0.695 0.757 0.668 1.2 0.6 0.4 2.2 5.35 6.4
YOLO11s 0.819 0.758 0.838 0.742 1.2 1 0.4 2.6 18.4 21.4
YOLO11m 0.898 0.826 0.893 0.795 1.2 2.4 0.4 4 38.8 67.9
YOLO11l 0.862 0.839 0.889 0.794 1.2 3 0.4 4.6 49 86.8
YOLO11x 0.819 0.816 0.885 0.784 0.9 6.1 0.4 7.4 109 194.8
YOLOv12n 0.883 0.684 0.802 0.708 1.2 4.6 0.4 6.2 5.34 6.4
YOLOv12s 0.761 0.839 0.877 0.78 1.6 4.3 0.5 6.4 18.1 21.4
YOLOv12m 0.868 0.811 0.89 0.79 0.9 4.9 0.4 6.2 39 67.3
YOLOv12l 0.826 0.836 0.877 0.781 0.9 5.5 0.4 6.8 51.2 89.6
YOLOv12x 0.832 0.849 0.889 0.793 1 9.5 0.5 11 114 198.9
The results in Table VII showcase the performance of the YOLO models on the Africa Wildlife dataset. This dataset contains large object sizes focusing on the ability of YOLO models to predict large objects and their risk of overfitting due to the size of the dataset. The models demonstrate robust accuracy across the board, with the highest-performing models achieving a mAP50-95 ranging from 0.832 to 0.725. This relatively shorter range reflects the effectiveness of the models in detecting and classifying large wildlife objects by maintaining high accuracy.
Figure 14: mAP50 and mAP50-95 YOLO results on Africa wildlife dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
Figure 15: Precision vs. Recall based on size results on Africa wildlife dataset, with larger circles indicating larger model sizes.
Figure 16: Total processing time and GFLOPs count results on Africa wildlife dataset.
Accuracy:
As illustrated in Figure 14, YOLOv9s demonstrates exceptional performance with a high mAP50-95 of 0.832 and a mAP50 of 0.956, showcasing its robust accuracy across various IoU thresholds. YOLOv9c and YOLOv9t follow closely, with mAP50 scores of 0.96 and 0.948 and mAP50-95 scores of 0.83 and 0.825, respectively. These results highlight the YOLOv9 family’s ability to effectively learn patterns from a small sample of images, making it particularly suited for smaller datasets. In contrast, YOLOv5un, YOLOv10n, and YOLOv3u-tiny show lower mAP50-95 scores of 0.791, 0.786, and 0.725, indicating their limitations in accuracy. The underperformance of larger models like YOLO11x, YOLOv5ux, YOLOv5ul, and YOLOv10l can be attributed to overfitting, especially given the small dataset size.
Precision and Recall:
Figure 15 reveals that YOLO8l and YOLO11l achieve the highest precision and recall, with values of 0.942 and 0.937 for precision, and 0.898 and 0.896 for recall, respectively. Notably, YOLOv8n achieves similar results (0.932 for precision, 0.908 for recall) with a compact size of 6.55Mb, demonstrating its efficiency. In contrast, YOLOv3u and YOLOv5ul exhibit lower precision and recall scores (0.91 and 0.88 for YOLOv3u, 0.916 and 0.881 for YOLOv5ul), despite their larger sizes (204.86Mb for YOLOv3u, 106.85Mb for YOLOv5ul), which may be attributed to overfitting issues.
Computational Efficiency:
YOLOv10n, YOLOv8n, and YOLOv3u-tiny are the fastest models, achieving processing times of 2ms and 1.8ms, with GFLOPs counts of 8.2 and 19.1, respectively. The first two models share the same processing speed and GFLOPs count, as showcased in Figure 16. Conversely, YOLOv12x exhibits the slowest processing time at 12.8ms and a GFLOPs count of 198.5, followed by YOLOv9e at 8.5ms and 189.3 GFLOPs count. These results indicate that larger models tend to require more processing time and hardware usage compared to smaller models, emphasizing the trade-off between model size and processing efficiency.
Overall Performance:
YOLOv9t and YOLOv9s consistently excel across all metrics, delivering high accuracy while maintaining small model sizes, low GFLOPs, and short inference times, as shown in Table VII, and Figures 15 and 16. This demonstrates the robustness of YOLOv9’s smaller models and their effectiveness on small datasets. In contrast, YOLOv12x and YOLO5ux show suboptimal accuracy despite their larger sizes and longer inference times, likely due to overfitting. Most large models underperformed on this dataset, with the exception of YOLOv10x, which benefited from a modern architecture that prevents overfitting.
TABLE VIII: Evaluation results for the ships and vessels dataset.
Versions Precision Recall mAP50 mAP50-95
Preprocess
Time
Inference
Time
Postprocess
Time
Total
Time
Size GFLOPs
YOLOv3u 0.679 0.534 0.519 0.298 0.8 6.2 0.3 7.3 207.86 282.5
YOLOV3u tiny 0.647 0.511 0.489 0.273 1 0.7 0.3 2 24.44 18.9
YOLOv5un 0.635 0.532 0.514 0.298 1.5 0.6 0.3 2.4 5.65 7.2
YOLOv5us 0.653 0.541 0.518 0.299 1.2 0.8 0.3 2.3 18.58 24
YOLOv5um 0.667 0.541 0.526 0.308 0.9 2.1 0.3 3.3 50.54 64
YOLOv5ul 0.654 0.545 0.525 0.305 0.9 3.3 0.3 4.5 106.85 134.8
YOLOv5ux 0.668 0.555 0.531 0.309 0.8 6.7 0.3 7.8 195.2 246.2
YOLOv8n 0.655 0.533 0.515 0.297 1.5 0.5 0.3 2.3 6.55 8
YOLOv8s 0.647 0.545 0.518 0.301 1.1 1 0.3 2.4 22.59 28.5
YOLOv8m 0.669 0.547 0.525 0.302 0.8 2.5 0.3 3.6 52.12 79
YOLOv8l 0.659 0.551 0.526 0.303 0.9 3.9 0.3 5.1 87.77 165
YOLOv8x 0.655 0.55 0.529 0.306 0.8 7.1 0.3 8.2 136.9 257.7
YOLOv9t 0.647 0.516 0.512 0.3 1.4 1.1 0.3 2.8 4.93 7.5
YOLOv9s 0.655 0.552 0.522 0.308 1.4 1.2 0.3 2.9 15.33 26.9
YOLOv9m 0.668 0.551 0.529 0.307 1.1 2.7 0.3 4.1 40.98 76.8
YOLOv9c 0.663 0.547 0.523 0.303 1.2 3.4 0.3 4.9 51.8 102.4
YOLOv9e 0.667 0.537 0.524 0.308 1.1 7.6 0.3 9 117.5 189.5
YOLOv10n 0.584 0.487 0.506 0.31 1.4 0.8 0.2 2.4 5.59 8.2
YOLOv10s 0.586 0.511 0.515 0.319 1.1 1.1 0.2 2.4 15.9 24.4
YOLOv10m 0.588 0.517 0.522 0.322 1 2.4 0.1 3.5 32.1 63.4
YOLOv10b 0.603 0.509 0.523 0.319 1.1 3.2 0.1 4.4 39.7 97.9
YOLOv10l 0.601 0.511 0.522 0.322 1.1 3.8 0.1 5 50 126.3
YOLOv10x 0.6 0.523 0.526 0.321 1 6.3 0.2 7.5 61.4 169.8
YOLO11n 0.574 0.51 0.505 0.311 1.5 0.7 0.3 2.5 5.35 6.3
YOLO11s 0.585 0.535 0.521 0.323 1.3 1 0.3 2.6 18.4 21.3
YOLO11m 0.588 0.541 0.53 0.325 1 2.4 0.3 3.7 38.8 67.6
YOLO11l 0.596 0.531 0.528 0.325 1.1 3 0.4 4.5 49 86.6
YOLO11x 0.596 0.538 0.529 0.327 0.8 6.1 0.3 7.2 109 194.4
YOLOv12n 0.571 0.505 0.505 0.307 1.1 4.7 0.4 6.2 5.34 6.3
YOLOv12s 0.58 0.529 0.516 0.319 1.6 5 0.4 7 18.1 21.2
YOLOv12m 0.584 0.543 0.525 0.325 1.1 4.6 0.4 6.1 39 67.1
YOLOv12l 0.588 0.532 0.522 0.323 2.6 5.2 0.4 8.2 51.2 88.5
YOLOv12x 0.586 0.538 0.52 0.321 1.2 9.3 0.4 10.9 114 198.5
Figure 17: mAP50 and mAP50-95 YOLO results on ships and vessel dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
Figure 18: Precision vs. Recall based on size results on ships and vessels dataset, with larger circles indicating larger model sizes.
Figure 19: Total processing time and GFLOPs count results on ships and vessels dataset.
IV-A3 Ships and Vessels Dataset
Table VIII presents the performance of YOLO models on the Ships and Vessels dataset, a large dataset featuring tiny objects with varying orientations. Overall, the models demonstrated moderate effectiveness in detecting ships and vessels, with mAP50-95 ranging from 0.273 to 0.327. This performance suggests that YOLO algorithms may face challenges in accurately detecting smaller objects with different orientations.
Accuracy:
The disparity between mAP50-95 and mAP50, illustrated in Figure 17, underscores the challenges YOLO models face with higher IoU thresholds when detecting small objects. Additionally, YOLO models struggle with detecting objects of varying rotations. Among the models, YOLO11x achieved the highest accuracy, with a mAP50 of 0.529 and a mAP50-95 of 0.327, closely followed by YOLO11l, YOLO11m, and YOLO11s, which recorded mAP50 values of 0.529, 0.528, and 0.53, and mAP50-95 values of 0.327, 0.325, and 0.325, respectively. These results highlight the robustness of the YOLO11 family in detecting small and tiny objects. In contrast, YOLOv3u-tiny, YOLOv8n, YOLOv3u, and YOLOv5n exhibited the lowest accuracy, with mAP50 scores of 0.489, 0.515, 0.519, and 0.514, and mAP50-95 scores of 0.273, 0.297, 0.298, and 0.298, respectively. This suggests the outdated YOLOv3u architecture and the potential underfitting of smaller models with large size of datasets.
Precision and Recall:
Figure 18 indicates that YOLOv5ux outperformed other models, achieving a precision of 0.668 and a recall of 0.555. It was closely followed by YOLOv9m (precision of 0.668, recall of 0.551) and YOLOv8m (precision of 0.669, recall of 0.525), both of which are significantly smaller in size (40.98 Mb for YOLOv9m and 52.12 Mb for YOLOv8m). In contrast, YOLOv12n and YOLO11n exhibited lower performance, with precisions of 0.571 and 0.574 and recalls of 0.505 and 0.51, respectively, likely due to underfitting issues. Generally, YOLOv12 and YOLO11 models tended to produce false positives, reflected in their low precision and high recall. Meanwhile, YOLOv10 underperformed in both precision and recall, despite being one of the newest models in the YOLO family.
Computational Efficiency:
As illustrated in Figure 19, YOLOv3u-tiny achieved the fastest processing time at 2 ms, closely followed by YOLOv8n and YOLOv5un, both recording 2.3 ms. YOLOv10 and YOLO11 models also excelled in speed, with YOLOv10n and YOLO11n achieving rapid inference times of 2.4 ms and 2.5 ms, along with GFLOPs counts of 8.2 and 6.3, respectively. In contrast, YOLOv12x exhibited the lowest speed, with 10.9 ms inference time and 198.5 GFLOPs count, highlighting the shortcomings of the newest YOLO family.
Overall Performance:
The results in Table VIII and Figures 17, 18, and 19 demonstrate that YOLO11s and YOLOv10s excelled in accuracy while maintaining compact sizes, low GFLOPs, and quick processing times. In contrast, YOLOv3u, YOLOv8x, and YOLOv8l fell short of expectations despite their larger sizes and longer processing times. These findings highlight the robustness and reliability of the YOLO11 family, particularly in improving the YOLO family’s performance in detecting small and tiny objects while ensuring efficient processing. Additionally, the results reveal the underperformance of YOLOv9 models when faced with large datasets and small objects, despite their modern architecture.
IV-B Discussion
Based on the performance of the models across the three datasets, we ranked them by accuracy, speed, GFLOps count, and size, as shown in Table IX to facilitate a comprehensive evaluation. For accuracy, the mAP50-95 metric was employed due to its capacity to assess models across a range of IoU thresholds, thus providing a detailed insight into each model’s performance. For speed, models were sorted based on the total processing time, which encompasses preprocessing, inference, and postprocessing durations. The rankings range from Rank 1, indicating the highest performance, to Rank 28, denoting the lowest, with the respective rankings highlighted in bold within the table.
TABLE IX: Overall ranking of YOLO algorithms
Version Acc. Speed GFLOPs Size
YOLOv3u tiny 33 1 7 33
YOLOv3u 25 28 33 13
YOLOv5un 32 6 3 5
YOLOv5us 29 8 10 11
YOLOv5um 21 15 15 21
YOLOv5ul 18 21 25 27
YOLOv5ux 22 31 31 32
YOLOv8n 31 5 5 6
YOLOv8s 26 9 13 12
YOLOv8m 19 18 19 24
YOLOv8l 17 23 26 26
YOLOv8x 11 30 32 31
YOLOv9t 23 14 4 1
YOLOv9s 10 16 12 7
YOLOv9m 5 22 18 18
YOLOv9c 13 26 23 23
YOLOv9e 16 33 28 30
YOLOv10n 30 2 6 4
YOLOv10s 23 3 11 8
YOLOv10m 9 9 14 14
YOLOv10b 11 12 22 17
YOLOv10l 14 16 24 20
YOLOv10x 2 26 27 25
YOLO11n 27 3 1 3
YOLO11s 19 6 8 10
YOLO11m 1 11 17 15
YOLO11l 2 13 20 19
YOLO11x 8 23 29 28
YOLOv12n 27 19 1 2
YOLOv12s 7 25 8 9
YOLOv12m 5 20 16 16
YOLOv12l 4 28 21 22
YOLOv12x 15 32 30 29
The analysis of Table IX yields critical observations:
IV-B1 Accuracy
YOLO11m consistently emerged as a top performer, frequently ranking among the highest, closely followed by YOLOv10x, YOLO11l, YOLOv9m, and YOLO11x. This underscores the robust performance of the YOLO11 family across varying IoU thresholds and object sizes, which can be attributed to their use of C2PSA for the preservation of contextual information, leading to improved convergence and overall performance. The inclusion of two small kernel convolutions and partial self-attention modules in these models helped capture more spatial features, leading to better accuracy in detecting both large and small objects. Additionally, the fine-tuning of anchor boxes and multi-scale training has enhanced the detection of objects at different resolutions, which is crucial in achieving high accuracy across diverse datasets.
Conversely, YOLOv3u-tiny exhibited the lowest accuracy, particularly in the Africa Wildlife and Ships and Vessels datasets. YOLOv5un and YOLOv8n show slightly better but still sub-par results. This suggests that YOLO11 models are currently the most reliable for applications demanding high accuracy.
The YOLOv3u-tiny model’s underperformance can be attributed to its older architecture, which lacks the optimizations found in newer YOLO models. YOLOv3u is based on the Darknet-53 framework, which, although efficient for its time, struggles to capture complex patterns and fine details required for high-accuracy detection. Its relatively shallow architecture and absence of more recent enhancements such as attention mechanisms or advanced feature fusion techniques make it less effective in handling intricate object detection tasks, especially in datasets with a wide variety of object sizes, like those in the Africa Wildlife and Ships and Vessels datasets.
In contrast, the YOLOv9 models show effectiveness in detecting objects across various sizes and IoU thresholds, but they still struggle with small objects, particularly in the Ships and Vessels dataset. The YOLOv9 family benefits from improved training strategies, such as anchor box refinement and better scale handling, but may still fall short in detecting small objects in crowded or cluttered scenes due to its reliance on larger-scale feature maps and a less efficient processing pipeline for fine-grained features.
The YOLOv10 family, despite its later introduction, exhibited relatively lower accuracy in the Traffic Signs and Africa Animals datasets, resulting in an average accuracy drop of 2.075% compared to the YOLOv9 models in those datasets. YOLOv10’s adoption of the One-to-One Head approach instead of Non-Maximum Suppression (NMS) for defining bounding boxes contributed to this decrease. The One-to-One Head approach relies on predicting a single bounding box per object, which can be a limitation in scenarios with overlapping or closely packed objects. NMS, on the other hand, is more effective at filtering out redundant predictions, which can improve detection accuracy in these types of datasets. Therefore, the One-to-One Head’s inability to handle multiple overlapping predictions adequately explains YOLOv10’s underperformance, especially in datasets like Traffic Signs, where multiple signs are often close together.
Although the YOLOv12 family includes strong models like YOLOv12l, which ranked 4th in accuracy, it falls short compared to previous iterations, particularly considering its status as the newest YOLO version. It shows a -2.72% decline compared to YOLOv9 in the African Animals dataset and a -1% decrease in accuracy compared to YOLO11 in the Ships and Vessels dataset. This decline in accuracy is largely due to architectural choices. R-ELAN and area attention, while innovative, compromised global dependency modeling, impacting accuracy. Additionally, the adjusted MLP ratio prioritized certain computational aspects over feature representation depth, limiting its ability to capture intricate patterns. Consequently, YOLOv12 underperformed compared to the more balanced YOLOv11 and YOLOv9 families, demonstrating that newer architectural changes did not yield the expected gains in accuracy.
Similarly, YOLOv8n, though it performed better than YOLOv3u, still showed lower accuracy than the top-performing models in specific datasets. This model demonstrated a good balance between speed and accuracy but struggled with intricate object recognition tasks. The simplicity of its architecture, without more advanced modules like self-attention or multi-scale training, likely held it back from achieving the high accuracy observed in models like YOLO11 and YOLOv9.
Lastly, the YOLO11 models stand out for their ability to consistently achieve the highest accuracy. The use of advanced mechanisms like C2PSA and two kernel convolutions allowed for improved feature extraction across multiple scales, enhancing their ability to detect objects of varying sizes and complexities. These models’ robust performance across a variety of datasets, including challenging ones like Africa Wildlife and Ships and Vessels, demonstrates their ability to handle a wide range of object detection tasks with exceptional accuracy.
IV-B2 Computational Efficiency
YOLOv10n consistently outperformed other models in terms of speed and GFLOPs count, ranking among the top for speed across all datasets and 5th in terms of GFLOPs. Its use of CSPNet as backbone design and optimization strategies allowed for high processing speed without significantly compromising performance. YOLOv3u-tiny, YOLOv10s, and YOLO11n also demonstrated excellent computational efficiency, with their lightweight architectures enabling faster inference times and lower resource usage.
Despite its accuracy advantages, YOLOv12 ranked as the worst family in terms of processing time and second in GFLOPs count. This is largely due to its attention-centric architecture, which enhances accuracy but significantly increases computational complexity. The integration of Area Attention and R-ELAN, while effective in capturing spatial features, introduces substantial overhead in feature extraction and aggregation, leading to slower inference times. Additionally, the reliance on FlashAttention contributes to inefficient memory access patterns, further impacting processing speed. Although these architectural choices improve contextual feature extraction, they hinder computational efficiency, making YOLOv12 less suitable for real-time applications compared to its predecessors. This performance trade-off illustrates that while attention mechanisms can elevate accuracy, they come at the cost of increased latency and computational demands.
YOLOv5un, YOLOv5us, and YOLOv8n exhibited strong efficiency, providing a good balance between speed and accuracy. These models performed well in both inference times and GFLOPs, with YOLOv5un showing slightly better speed than YOLOv8n potentially due to the implementation of C2k (CSP with larger kernel sizes) in YOLOv8 architecture that increases accuracy but with a trade-off in computational efficiency. YOLOv8n, despite its good speed performance, had a larger size and computational cost compared to some other lightweight models.
While YOLOv3u-tiny ranked highly in speed, its performance in accuracy was subpar, particularly in datasets like Africa Wildlife and Ships and Vessels. This highlights a critical trade-off: while speed is crucial for real-time applications, it becomes irrelevant if it severely compromises accuracy. YOLOv3u-tiny’s lightweight architecture, while optimized for faster inference, failed to detect objects accurately across varied and complex datasets, making it unsuitable for applications where precision is key. On the other hand, YOLOv3u struggled in both speed and accuracy, ranking poorly in both categories. Its reliance on the outdated Darknet-53 backbone contributed to this underperformance, as it was unable to meet modern detection challenges. The model’s low ranking in both computational efficiency and accuracy makes it less viable compared to newer architectures, such as YOLOv11 and YOLOv10, which offer better performance across both metrics.
In contrast, YOLO11 models demonstrated significant speed improvements, thanks to the C3k2 block, which reduces computational overhead while maintaining high accuracy. These models were faster than YOLOv10 by 1.41% and YOLOv9 by 31% on average, making them ideal for real-time applications requiring rapid processing without sacrificing accuracy. While YOLOv9 models excelled in accuracy, their slower inference times made them less ideal for time-sensitive tasks. YOLOv10, slightly slower than YOLO11, offered a good balance between speed and accuracy, making it well-suited for real-time applications.
Lastly, YOLOv5 and YOLOv8 families also performed well, with YOLOv5un ranking near the top for speed and efficiency, making it a solid choice for applications where both quick processing and solid accuracy are required.
IV-B3 Model Size
YOLOv9t emerged as the smallest model, ranking first across all three datasets in terms of model size thanks to its use of GELAN architecture leading to more efficient use of parameters. It is followed closely by YOLOv12n, YOLO11n, and YOLOv10n. Specifically, YOLOv10n benefits from the Spatial-Channel Decoupled Downsampling technique, which effectively reduces the model size while maintaining high accuracy, making it a highly efficient choice for resource-constrained environments. This technique allows for better utilization of parameters, enabling smaller models to perform similarly to larger counterparts.
In contrast, YOLOv3u was the largest model, ranking poorly in terms of both size and accuracy. This inefficiency stems from its reliance on the outdated Darknet-53 architecture, which is larger and more computationally expensive compared to more modern designs. The size of YOLOv3u further exacerbates its subpar performance, as the model requires more resources and longer inference times, without corresponding improvements in detection accuracy. Its larger size makes it less suitable for deployment in environments where computational resources are limited or where real-time processing is essential.
This comparison highlights the trend in newer YOLO versions, such as YOLOv9, YOLO10, and YOLO11, where optimizations have led to smaller, more efficient models without significant sacrifices in accuracy or speed. These advancements make them more attractive for modern applications, where both model size and computational efficiency are critical. However, despite being the latest addition, YOLOv12 did not follow this trend, ranking third in terms of model size and computational efficiency.
IV-B4 Overall Performance
Considering accuracy, speed, size, and GFLOPs, YOLO11m, YOLO11n, YOLO11s, and YOLOv10s emerged as the most consistent performers. These models demonstrated a strong balance between high accuracy, low processing time, and efficient disk usage, making them suitable for a wide range of applications where both speed and accuracy are crucial. YOLO11 models, in particular, benefited from recent enhancements such as the C3k2 block and C2PSA module, which contributed to their superior performance in these areas. The incorporation of these innovations allowed the YOLO11 models to handle larger datasets more effectively without compromising on speed or efficiency.
On the other hand, YOLOv9e, YOLOv5ux, and YOLOv3u showed poor results across all metrics. These models were computationally inefficient, struggling with high GFLOPs counts and slower processing times. YOLOv3u, in particular, ranked poorly due to its reliance on the outdated Darknet-53 architecture, which is no longer optimized for modern object detection tasks. Despite its larger size, it failed to offer the necessary accuracy, making it less suitable for practical applications. Similarly, YOLOv5ux and YOLOv9e showed poor performance in terms of both speed and accuracy, which further limits their real-world usability.
YOLO11 models were clearly the best performers overall, owing to their modern design and architectural improvements, making them highly suitable for a wide range of use cases. YOLOv10 models, although slightly underperforming in accuracy, excelled in computational efficiency due to their use of the One-to-One head for predictions. This design allowed YOLOv10 to process data more quickly and efficiently, making it ideal for time-sensitive applications where speed is prioritized over perfect accuracy.
While YOLOv9 models showed underperformance in terms of computational efficiency, they still remain competitive with YOLOv10 and YOLO11 in terms of accuracy, particularly due to their PGI integration, which helps in enhancing detection precision. This makes YOLOv9 a viable option for applications where high accuracy is prioritized over speed.
Concerning the YOLOv12 family, it showcased mediocre results, underperforming compared to its predecessors, particularly YOLOv9 and YOLO11, in both accuracy and speed. Despite introducing innovative components like the Area Attention Module (A2), Residual Efficient Layer Aggregation Networks (R-ELAN), and FlashAttention, these architectural changes contributed to increased computational overhead and slower inference times. The Area Attention Module segmented the feature map to maintain a large receptive field but limited contextual understanding, leading to lower accuracy. Meanwhile, R-ELAN improved feature aggregation but introduced latency due to its complex residual connections. FlashAttention aimed to optimize memory access but added processing overhead, further slowing down the model. Although YOLOv12 leveraged convolution operators for efficiency, it failed to offset the complexity introduced by its attention mechanisms. Overall, YOLOv12 struggled to balance speed and accuracy, making it less favorable compared to the more efficient and accurate YOLOv9, YOLOv10, and YOLO11 models.
In addition, YOLOv8 and YOLOv5u achieved competitive results, surpassing YOLOv3u in terms of accuracy. This improvement can be attributed to their more recent architectures, which allowed them to handle modern detection tasks better. However, despite their higher accuracy compared to YOLOv3u, they still fell short when compared to newer models like YOLOv9, YOLOv10, and YOLO11 in both accuracy and overall performance. While YOLOv8 and YOLOv5u had faster processing times than YOLOv9, their overall performance was still inferior to that of the YOLO11 and YOLOv10 models, which maintained a better balance between speed and accuracy.
IV-B5 The Rise of YOLO11 Over YOLOv8
Although YOLOv8 [27] has been the algorithm of choice for its versatility in tasks such as pose estimation, instance segmentation, and oriented object detection (OBB), YOLO11 [28] has now emerged as a more efficient and accurate alternative. With its ability to handle the same tasks while offering improved contextual understanding and better architectural modules, YOLO11 sets a new standard in performance, surpassing YOLOv8 in both speed and accuracy across various applications.
This paper contributes by providing empirical evidence of YOLO11’s superior performance in both accuracy and computational efficiency. Our findings establish YOLO11 as a clear advancement in the YOLO family, setting a new benchmark for object detection. By incorporating novel modules, YOLO11 addresses both the need for high accuracy and low-latency processing, offering valuable insights for deploying YOLO-based models in practical, real-world applications.
IV-B6 YOLOv9: Efficient Yet Slow
Although YOLOv9 [70] excels in GFLOPs count and model size, outperforming many other models, it paradoxically ranks as one of the slowest in terms of inference speed. This unusual combination of high computational efficiency but low speed can be attributed to YOLOv9’s architectural choices, such as the use of Programmable Gradient Information (PGI) and Generalized Efficient Layer Aggregation Network (GELAN), which optimize computational performance but add significant latency during inference.
This paper contributes by highlighting the trade-offs in YOLOv9’s design, showing that while its smaller model size and lower GFLOPs count make it ideal for resource-constrained environments, its slow inference times make it unsuitable for time-sensitive applications. Our findings suggest that YOLOv9’s performance is best suited for scenarios where computational efficiency is prioritized over speed, offering valuable insights for selecting models based on specific application needs.
IV-B7 YOLOv12: Ambitious but Underwhelming
YOLOv12, despite being the latest addition to the YOLO family, delivered subpar results, failing to outperform its predecessors, particularly YOLOv9, YOLOv10, and YOLO11, in both accuracy and speed. This underperformance is primarily due to its complex architectural choices, including the Area Attention Module (A2), Residual Efficient Layer Aggregation Networks (R-ELAN), and FlashAttention, which introduced significant computational overhead. Although these modules aimed to enhance contextual understanding and memory efficiency, they ultimately compromised speed and accuracy due to increased latency and reduced spatial awareness. This work contributes by revealing the trade-offs inherent in YOLOv12’s ambitious design, demonstrating that its added complexity did not translate into superior performance.
IV-B8 Dataset Size
The size of the dataset significantly influences the performance of YOLO models. For instance, large models did not perform optimally on the small African wildlife dataset compared to their results on the Traffic Signs and Ships and Vessels datasets due to being more prone to overfitting. Conversely, small models like YOLOv9t and YOLOv9s performed best on the Africa Wildlife dataset, showcasing the effectiveness of small-scaled models when handling limited datasets.
IV-B9 Impact of Training Datasets
The performance of YOLO models is influenced by the training datasets used, as shown in Tables VI, VII, and VIII. Different datasets yield varying results and top performers, indicating that dataset complexity affects algorithm performance. This underscores the importance of using diverse datasets during benchmarking to obtain comprehensive results on the strengths and limitations of each model.
This discussion highlights the need for a balanced consideration of accuracy, speed, and model size when selecting YOLO models for specific applications. The consistent performance of YOLO11 models across various metrics makes them highly recommended for versatile situations where accuracy and speed are essential. Meanwhile, YOLOv10 models can perform similarly while achieving faster processing times and with smaller model sizes. Additionally, YOLOv9 can deliver comparable results in terms of accuracy but sacrifices speed, making it suitable for applications where precision is prioritized over rapid processing.
V Conclusion
This benchmark study thoroughly evaluates the performance of various YOLO algorithms. It pioneers a comprehensive comparison of YOLO11 against its predecessors, evaluating their performance across three diverse datasets: Traffic Signs, African Wildlife, and Ships and Vessels. The datasets were carefully selected to encompass a wide range of object properties, including varying object sizes, aspect ratios, and object densities. We showcase the strengths and weaknesses of each YOLO version and family by examining a wide range of metrics such as Precision, Recall, Mean Average Precision (mAP), Processing Time, GFLOPs count, and Model Size. Our study addresses the following key research questions:
  • Which YOLO algorithm demonstrates superior performance across a comprehensive set of metrics?
  • How do different YOLO versions perform on datasets with diverse object characteristics?
  • What are the specific strengths and limitations of each YOLO version, and how can these insights inform the selection of suitable algorithms for various applications?
In particular, the YOLO11 family emerged as the most consistent, with YOLO11m striking an optimal balance between accuracy, efficiency, and model size. While YOLOv10 delivered slightly lower accuracy than YOLO11, it excelled in speed and efficiency, making it a strong choice for applications requiring efficiency and fast processing. Additionally, YOLOv9 performed well overall and particularly stood out in smaller datasets. These findings provide valuable insights for industry and academia, guiding the selection of the most suitable YOLO algorithms and informing future developments and enhancements.
YOLOv12, despite being the newest addition to the YOLO family, delivered underwhelming performance. Its ambitious architectural changes, including the Area Attention Module and R-ELAN, introduced complexity without translating into clear advantages. The model struggled to balance accuracy and speed, highlighting the challenges of integrating advanced attention mechanisms effectively within the YOLO framework.
While the evaluated algorithms show promising performance, there is room for improvement. Future research should focus on optimizing YOLOv10’s accuracy while maintaining its speed and efficiency. Additionally, simplifying YOLOv12’s complex attention mechanisms could enhance its speed without compromising accuracy. Further architectural advancements could lead to even more efficient and accurate YOLO models. Our future work will address these gaps and explore enhancements to maximize overall efficiency and real-world applicability.

### v4

@@
-YOLO Evolution: A Comprehensive Benchmark and Architectural Review of YOLOv12, YOLO11, and Their Previous Versions
+
 Abstract
@@
 Our analysis highlights the distinctive strengths and limitations of each YOLO version.
-For example: YOLOv9 demonstrates substantial accuracy but struggles with detecting small objects and efficiency whereas YOLOv10 exhibits relatively lower accuracy due to architectural choices that affect its performance in overlapping object detection but excels in speed and efficiency.
+For example, YOLOv9 demonstrates substantial accuracy but struggles with detecting small objects and efficiency, whereas YOLOv10 exhibits relatively lower accuracy due to architectural choices that affect its performance in overlapping object detection but excels in speed and efficiency.
 Additionally, the YOLO11 family consistently shows superior performance maintaining a remarkable balance of accuracy and efficiency.
@@
 These results provide critical insights for both industry and academia, facilitating the selection of the most suitable YOLO algorithm for diverse applications and guiding future enhancements.
-Index Terms:
-YOLO (You Only Look Once), YOLOv12, YOLO11, Object detection, Ultralytics, Benchmark Analysis.
-I Introduction
+keywords:
+YOLOv12 , YOLO11 , Benchmark , Evaluation , Architectural Review
+journal: Image and Vision Computing
+1 Introduction
 Object detection is an essential component of computer vision systems, enabling automated identification and localization of objects within images or video frames [44].
-Its applications span from autonomous driving and robotics [17, 5, 21] [59] to inventory management, video surveillance, and sports analysis [4, 24, 58, 74].
+Its applications span from autonomous driving and robotics [20, 7, 26, 69] to inventory management, video surveillance, and sports analysis [6, 30, 68, 87].
 Over the years, object detection has developed significantly.
@@
 In contrast, two-stage methods, such as Region-based Convolutional Neural Networks (R-CNN) [24], generate region proposals and then perform classification, offering high precision but being computationally intensive.
-Among one-stage object detection methods, YOLO (You Only Look Once) stands out for its robustness and efficiency.
-Initially introduced in 2015 by Redmon et al.
-[22], YOLO redefined object detection by predicting bounding boxes and class probabilities directly from full images in a single evaluation [50].
-This innovative approach allowed YOLOv1 to achieve real-time object detection with impressive accuracy.
-Building upon this foundation, YOLOv2 [51] incorporated several key enhancements.
-It integrated the Darknet-19 framework, a 19-layer convolutional neural network that improved feature extraction.
-YOLOv2 also introduced batch normalization and employed data augmentation techniques inspired by the VGG architecture [60] to enhance the model’s generalization.
-YOLOv3 [52] further advanced the model with the Darknet-53 framework, a deeper network that significantly improved feature extraction capabilities.
-This version also utilized a Feature Pyramid Network (FPN)-inspired design, which allowed for better detection across various object scales by combining high-level semantic features with low-level detailed features, and a Three-Scale detection mechanism that improved accuracy for objects of different sizes.
-Following YOLOv3, the model’s development branched into various communities, leading to several notable iterations.
-YOLOv4 [7], developed by Bochkovskiy et al., introduced enhancements such as Spatial Pyramid Pooling (SPP) and the Path Aggregation Network (PAN).
-SPP aggregates features from multiple scales, preserving spatial information, while PAN improves the fusion of features between layers, resulting in improved speed and accuracy.
-YOLOv5 [64] marked a significant transition by moving from the Darknet framework to PyTorch, a popular deep learning library.
-This transition made the model more accessible and easier to customize.
-The architecture incorporated strided convolution layers, which reduced computational load, and Spatial Pyramid Pooling Fast (SPPF) layers, optimizing memory usage while maintaining high performance.
-YOLOv6 and YOLOv7 continued this trajectory with innovative architectures.
-YOLOv6 [31] introduced RepVGG, an architecture that simplified convolutional layers during inference, and CSPStackRep blocks, which improve accuracy by splitting the feature map into two parts to process them separately.
-In addition, YOLOv6 employed a hybrid channel strategy for better feature representation.
-YOLOv7 [69] introduced the Extended Efficient Layer Aggregation Network (E-ELAN) to improve efficiency and effectiveness by enhancing information flow between layers.
-Figure 1: Evolution of YOLO Algorithms throughout the years.
-The most recent versions of YOLO, including YOLOv8, YOLOv9, YOLOv10, YOLO11, and YOLOv12 represent the forefront of the model’s development.
-YOLOv8 [61], released by Ultralytics, introduced semantic segmentation capabilities, allowing the model to classify each pixel of an image, and provided scalable versions to meet various application needs, from resource-constrained environments to high performance systems alongside other tasks such as pose estimation, image classification, and oriented object detection (OOB).
-YOLOv9 [70] built on its predecessors’ architectural advancements with Programmable Gradient Information (PGI), which optimizes gradient flow during training, and the Generalized Efficient Layer Aggregation Network (GELAN), which further improved performance by enhancing layer information flow.
-YOLOv10 [68], developed by Tsinghua University, eliminated the need for Non-Maximum Suppression (NMS) used by its predecessors, a technique used to eliminate duplicate predictions and pick the bounding boxes with the most confidence, by introducing a dual assignment strategy in its training protocol.
-Additionally, YOLOv10 features lightweight classification heads, spatial-channel decoupled downsampling, and rank-guided block design, making it one of the most efficient and effective YOLO models to date.
-YOLO11 [28], also introduced by Ultralytics, retains the capabilities of YOLOv8 with applications such as Instance Segmentation, Pose Estimation, and Oriented Object Detection while providing 5 scalable versions for different use cases.
-YOLO11 replaces the C2f block from YOLOv8 with the more efficient C3k2 block, delivering improved performance without compromising speed.
-Additionally, it introduces the C2PSA (Cross Stage Partial with Spatial Attention) module, which improves spatial attention in feature maps, increasing accuracy, especially for small and overlapping objects.
-Lastly, YOLOv12 joins the YOLO family by also implementing applications such as Pose Estimation, Instance Segmentation, and Oriented Object Detection.
-Additionally, it provides 5 scalable version and introduces new architectural advancements such as the Area Attention (A2) and the Residual Efficient Layer Aggregation Networks (R-ELAN), which improve training stability and model convergence through block-level residual design and optimized feature aggregation.
-This object detection algorithm has undergone several developments, as seen in Figure reffig:yoloevolution achieving competitive results in terms of accuracy and speed, making it the preferred algorithm in various fields such as ADAS (Advanced Driver-Assist System) [50], video surveillance [41], face detection [42], and many more [19].
-For instance, YOLO plays a crucial role in the agriculture field by being implemented in numerous applications such as crop classification [1] [18], pest detection [36], automated farming [71] [40], and virtual fencing [66].
-Moreover, YOLO has been utilized on numerous occasions in the field of healthcare such as cancer detection [48], ulcer detection [2], medicine classification [39] [45], and health protocols enforcement [12].
+YOLO (You Only Look Once) is a pioneering one-stage object detection framework that predicts bounding boxes and class probabilities in a single evaluation, enabling real-time performance.
+First introduced by Redmon et al.
+in 2015 [59], YOLOv1 redefined object detection with its efficiency.
+YOLOv2 [60] improved upon this by integrating Darknet-19, batch normalization, and data augmentation techniques inspired by the VGG architecture [70], leading to better generalization.
+YOLOv3 [61] adopted Darknet-53, a deeper network with enhanced feature extraction, and incorporated a Feature Pyramid Network (FPN)-inspired design for multi-scale detection, significantly improving accuracy for objects of varying sizes.
+Subsequent iterations diversified under different development groups.
+YOLOv4 [10] introduced Spatial Pyramid Pooling (SPP) for multi-scale feature fusion and the Path Aggregation Network (PAN) to refine feature integration.
+YOLOv5 [74] marked a shift from the Darknet framework to PyTorch, increasing accessibility and efficiency through strided convolution layers and Spatial Pyramid Pooling Fast (SPPF) layers.
+YOLOv6 [38] implemented RepVGG for simplified inference and CSPStackRep blocks for improved accuracy.
+YOLOv7 [81] introduced Extended Efficient Layer Aggregation Networks (E-ELAN) to optimize information flow and enhance detection performance.
+More recent versions further pushed the boundaries of efficiency and accuracy.
+YOLOv8 [71], released by Ultralytics, introduced scalable models tailored for various hardware constraints and expanded capabilities to tasks like semantic segmentation, pose estimation, and oriented bounding box (OBB) detection.
+YOLOv9 [82] introduced Programmable Gradient Information (PGI) for optimized gradient flow and Generalized Efficient Layer Aggregation Networks (GELAN) for enhanced feature fusion.
+YOLOv10 [79] eliminated Non-Maximum Suppression (NMS) using a dual assignment strategy, while also introducing lightweight classification heads and spatial-channel decoupled downsampling for increased efficiency.
+YOLOv11 [33] retained YOLOv8’s multi-task versatility while improving efficiency with the C3k2 block and introducing the C2PSA module for better spatial attention, particularly benefiting small and overlapping object detection.
+The latest iteration, YOLOv12, builds on prior advancements by incorporating an attention-centric approach with Area Attention (A2), which enhances feature aggregation while maintaining real-time performance.
+It also introduces Residual Efficient Layer Aggregation Networks (R-ELAN) to improve optimization stability and convergence [73].
+These innovations solidify YOLO as a leading framework in real-time object detection, continuously improving accuracy, speed, and adaptability across various tasks.
+Figure 1: Evolution of YOLO Algorithms throughout the years
+This object detection algorithm has undergone several developments, as seen in Figure 1 achieving competitive results in terms of accuracy and speed, making it the preferred algorithm in various fields such as ADAS (Advanced Driver-Assist System) [59], video surveillance [49], face detection [50], and many more [23].
+For instance, YOLO plays a crucial role in the agriculture field by being implemented in numerous applications such as crop classification [2] [22], pest detection [42], automated farming [85, 48], and virtual fencing [77].
+Moreover, YOLO has been utilized on numerous occasions in the field of healthcare, such as cancer detection [57], ulcer detection [3], medicine classification [47, 53], and health protocols enforcement [15].
 In recent years, Ultralytics has played a crucial role in the advancement of YOLO by maintaining, improving, and making these models more accessible [58].
-Notably, Ultralytics has streamlined the process of fine-tuning and customizing YOLO models, a task that was considerably more complex in earlier iterations.
+Notably, Ultralytics has streamlined the process of fine-tuning and customizing YOLO models, a considerably more complex task in earlier iterations.
 The introduction of user-friendly interfaces, comprehensive documentation, and pre-built modules has greatly simplified essential tasks such as data augmentation, model training, and evaluation.
 Moreover, the development of scalable model versions allows users to select models tailored to specific resource constraints and application requirements, thereby facilitating more effective fine-tuning.
-For instance, YOLOv8n is favorable over YOLOv8m in scenarios where speed and computational efficiency are prioritized over accuracy, making it ideal for resource-constrained environments.
 The integration of advanced tools for hyperparameter tuning, automated learning rate scheduling, and model pruning has further refined the customization process.
 Continuous updates and robust community support helped YOLO models to be more accessible and adaptable for a wide range of applications.
-This paper aims to present a comprehensive comparative analysis of the YOLO algorithm’s evolution.
+This paper presents a comprehensive comparative analysis of the YOLO algorithm’s evolution.
 It makes a significant contribution to the field by offering the first comprehensive evaluation of YOLOv12, the newest member of the YOLO family.
@@
 The rest of this paper is organized as follows: Section 2 covers related work.
-Section 3 describes the datasets, the models, and the experimental setup, including the hyperparameters and evaluation metrics used.
-Section 4 discusses the experimental results and comparative analysis.
+Section 3 describes the datasets, models, experimental setup, hyperparameters, and evaluation metrics used.
+Section 4 presents the experimental results, comparative analysis, and discussion, highlighting key discoveries, trends, architectural review, and real-life applications.
 Finally, Section 7 concludes with insights drawn from the study.
-II Related Work
+2 Related Work
 The YOLO (You Only Look Once) algorithm is considered one of the most prominent object detection algorithms.
@@
 The research aims to produce machine learning algorithms with cloud segmentation using a few spectral bands, including RGB and RGBN-IR combinations.
-Showcasing the implementation of YOLO in marine ship detection, the study [25] propose the Adaptive Multi-scale YOLO (AM YOLO) to tackle marine ship segmentation challenges, introducing modules for enhanced detail extraction and cross-scale integration.
-While achieving up to 4.0% precision gains over YOLOv5s, the study focuses primarily on marine environments, limiting generalization to other domains—a gap addressed in this paper through broader dataset evaluations and versatile benchmarks alongside implementing newer state-of-the-art YOLO models.
-In the paper [72], the authors developed a lightweight YOLO model optimized from YOLOv5, designed for multi-target garbage detection in dense and resource-constrained scenarios including MobileViTv3 for feature extraction, optimized MobileViT blocks for enhanced discrimination, and feature reuse techniques to minimize information loss.
-However, their reliance on YOLOv5 as a base architecture, which is less efficient and accurate than newer versions like YOLOv10 or YOLO11, limits both computational efficiency and performance—gaps addressed in this paper by leveraging state-of-the-art YOLO models and broader domain evaluations.
-In the field of Smart Cities, the authors [6] work on a comprehensive survey of 3D object detection methods in road scene understanding.
-Their systematization offers a practical framework for comparing approaches, aiding researchers and practitioners in navigating the field.
-However, the study primarily focuses on other object detection models neglecting the use of YOLO in 3D object detection.
-The paper [35] introduce the Efficient Attention Pyramid Transformer (EAPT), which enhances vision tasks by using Deformable Attention and the Encode-Decode Communication module (En-DeC) to capture more dynamic, global attention.
-While EAPT improves performance in various vision tasks, its integration with YOLO models is not explored.
-Incorporating these techniques into YOLO could enhance its feature extraction and global context understanding further improving the performance of YOLO variants.
-However, the study neglects the implementation of YOLO as a backbone, overlooking its potential in this area.
+Highlighting advancements in UAV-based object detection, the study [36] introduces SFFEF-YOLO to tackle challenges posed by varying target scales and small object prevalence in aerial imagery.
+By replacing the large prediction head with a tiny one, integrating a Fine-Grained Information Extraction Module (FIEM), and refining multi-scale fusion with the Multi-Scale Feature Fusion Module (MFFM), the model achieves an increase of 9.9% on VisDrone2019-DET and 3.6% on UAVDT in terms of mAP50 over YOLOv8.
+However, using an older YOLO version may limit its full potential.
+Incorporating newer YOLO models such as YOLO11 could further enhance accuracy, efficiency, and generalizability, a gap this paper addresses through broader benchmarks.
 The authors of [28] review the evolution of the YOLO variants from version 1 to version 8, examining their internal architecture, key innovations, and benchmarked performance metrics.
-The paper highlights the models’ applications across domains like autonomous driving and healthcare, and proposes incorporating federated learning to improve privacy, adaptability, and generalization in collaborative training.
+The paper highlights the models’ applications across domains like autonomous driving and healthcare and proposes incorporating federated learning to improve privacy, adaptability, and generalization in collaborative training.
 The review, however, limits its focus to mAP (mean Average Precision) for accuracy evaluation, neglecting other key metrics such as Recall and Precision.
@@
 However, the paper does not include a benchmarking study or a comparative analysis of the YOLO models, leaving out performance comparisons across the versions.
-The paper in [65] explores the development of the YOLO algorithm till the fourth version.
-It highlights its challenges and suggests new approaches, highlighting its impact on object detection and the need for ongoing study.
-However, it does not include newer YOLO models and lacks a comparative benchmark, limiting its ability to assess the progress and performance differences across various versions.
+The study [37] presents a lightweight rotational object detection algorithm to address challenges in remote sensing and surveillance, particularly variations in object size and orientation.
+By integrating an angle prediction branch and the Circular Smooth Label (CSL) method, along with a Channel and Spatial Attention (CSA) module, the model enhances feature extraction and detection accuracy.
+Achieving 57.86 mAP50 on the DOTA v2 dataset while maintaining a low computational footprint, it demonstrates efficiency for real-time deployment.
+However, its reliance on YOLOv5 limits potential performance gains.
+However, maintaining the backbone architecture of YOLOv5 can lead to limitations in terms of both computational efficiency and accuracy.
+Therefore, incorporating newer YOLO architectures could further improve the scope of this paper,
 The authors in the work in [34] analyze the YOLO algorithm, focusing on its development and performance.
@@
 The paper focuses exclusively on the mAP metric, overlooking other accuracy measures such as Precision and Recall.
-Additionally, it neglects speed and efficiency metrics limiting the scope of the comparative study.
-The paper also omits the evaluation of the most recent models, YOLOv9, YOLOv10, and YOLO11.
+Additionally, it neglects speed and efficiency metrics, limiting the scope of the comparative study.
+The paper also omits evaluating the most recent models, YOLOv9, YOLOv10, and YOLO11.
 This paper makes several key contributions: (i) It pioneers a comprehensive comparison of YOLOv12 against its predecessors across their scaled variants from nano- to extra-large; (ii) It provides deep insights into the architectural advancements of YOLO by analyzing key structural developments across versions.
@@
 (iv) The performance evaluation extends beyond mAP and FPS, incorporating critical metrics such as Precision, Recall, Preprocessing Time, Inference Time, Postprocessing Time, GFLOPs, and model size.
-(v) These comprehensive metrics provide valuable insights for selecting the most optimal YOLO models, benefiting both industry professionals and academics.
+(v) These comprehensive metrics provide valuable insights for selecting the most optimal YOLO models, benefiting industry professionals and academics.
 (vi) It offers specific use case recommendations, identifying the most suitable YOLO models for different scenarios and environments, such as resource-constrained deployments, real-time applications, and the detection of small or overlapping objects.
-III Benchmark Setup
-III-A Datasets
-This study aims to conduct in-depth benchmark research and assess the YOLO algorithms provided by the Ultralytics library.
+3 Benchmark Setup
+3.1 Datasets
+This study conducts in-depth benchmark research and assesses the YOLO algorithms provided by the Ultralytics library.
 The main goal is to provide a thorough and comparative analysis of these models and explain their strengths, deficiencies, and possible applications.
 This paper is made possible using several publicly accessible datasets on Kaggle and Roboflow.
-The selection of the datasets is based on the increasing implementation of the YOLO algorithms in the fields of Autonomous driving [50] [57] [32] [8], satellite imagery [33] [9] [47], and wildlife conservation [53] [73] [54].
+The selection of the datasets is based on the increasing implementation of the YOLO algorithms in the fields of Autonomous driving [59, 67, 39, 11], satellite imagery [40, 12, 56], and wildlife conservation [62, 86, 63].
 Moreover, each picked dataset presents unique difficulties and situations for object detection with varying image sizes and number of observations alongside the number of classes to ensure a comprehensive evaluation.
-III-A1 Traffic Signs Dataset
+3.1.1 Traffic Signs Dataset
 The Traffic Signs dataset by Radu Oprea is an open-source dataset on Kaggle containing around 55 classes across 3253 training and 1128 validation images of traffic signs in various sizes and environments [51].
@@
 To balance the classes, undersampling techniques were applied, as shown in Figure 2.
-After preprocessing the dataset by removing classes with low observation counts, 24 classes remained, with 3233 images split into 70% training, 20% validation, and 10% testing, with no data augmentation techniques applied.
+After preprocessing the dataset by removing classes with less than 50 observation counts, 24 classes remained, with 3233 images split into 70% training, 20% validation, and 10% testing, with no data augmentation techniques applied.
 This dataset is vital for applications in autonomous driving, traffic management, road safety, and intelligent transportation systems.
@@
 Figure 2: Classes Distribution of the Traffic Signs Dataset
-III-A2 Africa Wild Life Dataset
+3.1.2 Africa Wild Life Dataset
 The Africa Wildlife dataset is an open-source Kaggle dataset by Bianca Ferreira, designed for real-time animal detection in nature reserves [72].
@@
 The dataset presents challenges such as varying aspect ratios, with each image containing at least one instance of the specified animal class and potentially multiple occurrences of other classes.
-Furthermore, the overlapping of target objects increases the difficulty of detection.
+Furthermore, the overlap of target objects increases the difficulty of detection.
 This dataset is crucial for applications in wildlife conservation, anti-poaching efforts, biodiversity monitoring, and ecological research.
 Figure 3: Classes Distribution of the Africa Wildlife Dataset
-III-A3 Ships/Vessels Dataset
+3.1.3 Ships/Vessels Dataset
 The Ships/Vessels dataset is an extensive open-source collection containing approximately 13.5k images, collected by Siddharth Sah from numerous Roboflow datasets, curated explicitly for ship detection [65].
 Each image has been manually annotated with bounding boxes in the YOLO format, ensuring precise and efficient detection of ships.
-This dataset features a single class, ”ship,” and is divided into 70% training, 20% validation, and 10% testing allowing for streamlined and focused analysis.
+This dataset features a single class, ”ship,” and is divided into 70% training, 20% validation, and 10% testing.
 Notably, no data augmentation techniques were applied during the experiment, which ensures that the model’s performance is evaluated on the raw, unaltered dataset.
-However, the relatively small size of the target objects and their varying rotations pose challenges for detection, particularly for the YOLO algorithm, which often struggles with small object detection and objects with varying orientations.
+However, the relatively small size of the target objects, varying aspect ratios, and their varying rotations pose challenges for detection, particularly for the YOLO algorithm, which often struggles with small object detection and objects with varying orientations.
 The dataset is essential for various practical applications such as maritime safety, fisheries management, marine pollution monitoring, defense, maritime security, and more.
-III-B Models
-III-B1 Ultralytics vs. Original YOLO
-In this subsection, we will conduct a comparative analysis between the models provided by Ultralytics and their original counterparts on the Traffic Signs dataset provided by Radu Oprea [43] using the same hyperparameters in Table V.
+3.2 Models
+3.2.1 Ultralytics vs. Original YOLO
+In this subsection, we will conduct a comparative analysis between the models provided by Ultralytics and their original counterparts on the Traffic Signs dataset provided by Radu Oprea [51] using the same hyperparameters in Table 3.
 The objective is to highlight the differences between Ultralytics models and the original versions, which justifies the exclusion of YOLOv4 [10], YOLOv6 [38], and YOLOv7 [81] from this paper due to the lack of support for these models by Ultralytics.
@@
 Ultralytics Supported Models and Tasks:
-Ultralytics library provides researchers and programmers with various YOLO models for inference, validation, training, and export.
-Based on the results of Table I, we notice that Ultralytics does not support YOLOv1, YOLOv2, YOLOv4, and YOLOv7.
+Ultralytics library provides researchers and programmers various YOLO models for inference, validation, training, and export.
+Based on the results of Table 1, we notice that Ultralytics does not support YOLOv1, YOLOv2, YOLOv4, and YOLOv7.
 Concerning YOLOv6, the library only supports the configuration *.yaml files without the pre-trained *.pt models.
-TABLE I: Ultralytics-supported library tasks and models
+Table 1: Ultralytics-supported library tasks and models
 YOLO
@@
 00footnotetext: Boldface text shows available pre-trained models
+Figure 4: YOLO versions and scaled versions
 Performance Comparison of Ultralytics and Original Models:
-Based on the results of our comparative analysis of the Ultralytics models and their original counterparts on the Traffic Signs dataset presented in Table II, we observe significant discrepancies between the performance of the Ultralytics models and their original counterparts.
+Based on the results of our comparative analysis of the Ultralytics models and their original counterparts on the Traffic Signs dataset presented in Table 2, we observe significant discrepancies between the performance of the Ultralytics models and their original counterparts.
 Notably, Ultralytics’ versions of YOLOv5n (nano) and YOLOv3 demonstrate superior performance, underscoring the enhancements and optimizations implemented by Ultralytics.
@@
 Therefore, this paper will focus exclusively on the Ultralytics-supported versions to ensure consistent and fair benchmarks.
-In total, 33 models from 7 different YOLO versions were trained on 3 different datasets as seen in table III.
-Each YOLO version includes several scaled models, (e.g., YOLOv5u, YOLOv5un, YOLOv5us, and YOLOv5ux), with the suffixes denoting model size and complexity, such as ”n” for nano, ”s” for small, ”m” for medium, ”l” for large, ”x” for extra-large, ”t” for tiny, ”c” for compact, ”b” for balanced, and ”e” for extended.
+In total, 33 models from 7 different YOLO versions were trained on three different datasets, as seen in Figure 4.
+Each YOLO version includes several scaled models (e.g., YOLOv5u, YOLOv5un, YOLOv5us, and YOLOv5ux), with the suffixes denoting model size and complexity, such as ”n” for nano, ”s” for small, ”m” for medium, ”l” for large, ”x” for extra-large, ”t” for tiny, ”c” for compact, ”b” for balanced, and ”e” for extended.
 These models offer a variety of trade-offs between accuracy and inference speed, as discussed in the following sections.
-TABLE II: Ultralytics and original YOLO performance comparison
+Table 2: Ultralytics and original YOLO performance comparison
 Version Source mAP50 mAP50-95
@@
 Github 0.562 0.471
-TABLE III: YOLO versions and scaled versions
-Version Scaled Version
-YOLOv3u YOLOv3u-tiny
-YOLOv3u
-YOLOv5u YOLOv5un (nano)
-YOLOv5us (small)
-YOLOv5um (medium)
-YOLOv5ul (large)
-YOLOv5ux (extra-large)
-YOLOv8 YOLOv8n (nano)
-YOLOv8s (small)
-YOLOv8m (medium)
-YOLOv8l (large)
-YOLOv8x (extra-large)
-YOLOv9 YOLOv9t (tiny)
-YOLOv9s (small)
-YOLOv9m (medium)
-YOLOv9c (compact)
-YOLOv9e (extended)
-YOLOv10 YOLOv10n (nano)
-YOLOv10s (small)
-YOLOv10m (medium)
-YOLOv10b (balanced)
-YOLOv10l (large)
-YOLOv10x (extra-large)
-YOLO11 YOLO11n (nano)
-YOLO11s (small)
-YOLO11m (medium)
-YOLO11l (large)
-YOLO11x (extra-large)
-YOLOv12 YOLO11n (nano)
-YOLOv12s (small)
-YOLOv12m (medium)
-YOLOv12l (large)
-YOLOv12x (extra-large)
-III-B2 YOLOv3u
-YOLOv3, based on its predecessors, aims to improve localization errors and detection efficiency, particularly for smaller objects.
-It uses the Darknet-53 framework, which has 53 convolutional layers and achieves double the speed of ResNet-152 [52].
-YOLOv3 also incorporates elements from the Feature Pyramid Network (FPN), such as residual blocks, skip connections, and up-sampling, to enhance its ability to detect objects efficiently across varying scales, as seen in Figure 4.
-The algorithm generates feature maps at three distinct scales, down-sampling the input at factors of 32, 16, and 8, and uses a three-scale detection mechanism to detect large, medium, and small-sized objects using distinct feature maps.
-Despite its improvements, YOLOv3 still faces challenges in achieving precise results for medium and large-sized objects, so Ultralytics released YOLOv3u.
-YOLOv3u is an improved version of YOLOv3 that utilizes an anchor-free detection method used later in YOLOv8 and improves upon the accuracy and speed of YOLOv3, especially for medium and large-sized objects.
-Figure 4: YOLOv3 architecture showcasing the residual blocks and the upsampling layers [10].
-III-B3 YOLOv5u
-YOLOv5, proposed by Glenn Jocher, transitions from the Darknet framework to PyTorch, retaining many improvements from YOLOv4 [64] [26] and utilizing CSPDarknet as its backbone.
-CSPDarknet is a modified version of the original Darknet architecture that incorporates Cross-Stage Partial connections by splitting feature maps into separate paths, allowing for more efficient feature extraction and reduced computational costs.
-YOLOv5 features a strided convolution layer with a large window size, aiming to reduce memory and computational costs, as showcased in Figure 5.
-Moreover, this version adopts the Spatial Pyramid Pooling Fast (SPPF) module to provide a multiscale representation of the input feature maps.
-The SPPF module works by pooling features at different scales and concatenating them, allowing the network to capture fine and coarse information.
-This helps recognize objects of various sizes more effectively.
-In addition, YOLOv5 implements several augmentations, such as Mosaic, copy-paste, random affine, MixUp, HSV augmentation, and random horizontal flip.
-YOLOv5 is available in five variants, varying in width and depth of convolution modules.
-Ultralytics is actively improving this model through YOLOv5u, adopting an anchor-free detection method and achieving better overall performance, especially on complex objects of different sizes.
-Figure 5: Detailed architecture of YOLOv5 including the CSPDarknet Backbone, PANet Neck, and YOLO Layer Head [14].
-III-B4 YOLOv8
+3.2.2 YOLOv3u
+YOLOv3 enhances localization and detection efficiency, particularly for small objects, using the Darknet-53 framework, which offers double the speed of ResNet-152 [61].
+It integrates Feature Pyramid Network (FPN) elements, including residual blocks, skip connections, and up-sampling, to improve multi-scale detection, as illustrated in Figure 5.
+The model generates feature maps at three scales (down-sampling at factors of 32, 16, and 8) for detecting objects of varying sizes.
+However, YOLOv3 struggles with medium and large objects, leading Ultralytics to introduce YOLOv3u, which adopts an anchor-free detection method, later used in YOLOv8, improving both accuracy and speed.
+Figure 5: YOLOv3 architecture showcasing the residual blocks and the upsampling layers [13]
+3.2.3 YOLOv5u
+Figure 6: Detailed architecture of YOLOv5 including the CSPDarknet Backbone, PANet Neck, and YOLO Layer Head [17]
+YOLOv5, developed by Glenn Jocher, transitions from Darknet to PyTorch, incorporating CSPDarknet as its backbone for improved feature extraction and reduced computational cost [74] [31].
+It features Cross-Stage Partial (CSP) connections, a strided convolution layer, and the Spatial Pyramid Pooling Fast (SPPF) module, which enhances multi-scale feature representation, as illustrated in Figure 6.
+YOLOv5 also integrates data augmentation techniques (Mosaic, copy-paste, MixUp, HSV augmentation) to improve robustness.
+Available in five scaled variants, it continues to evolve with YOLOv5u, which introduces anchor-free detection for enhanced accuracy and efficiency, particularly on complex objects of varying sizes.
+3.2.4 YOLOv8
 Ultralytics has introduced YOLOv8, a significant evolution in the YOLO series, with five scaled versions [71] [32].
-Alongside object detection, YOLOv8 also provides various applications such as image classification, pose estimation, instance segmentation, and oriented object detection (OOB).
+Alongside object detection, YOLOv8 also provides various applications such as image classification, pose estimation, instance segmentation, and OBB.
 Key features include a backbone similar to YOLOv5, with adjustments in the CSPLayer, now known as the C2f module, which combines high-level features with contextual information for enhanced detection accuracy highlighted in Figure 7.
@@
 Figure 7: Detailed architecture of YOLOv8 showcasing the backbone’s multiple convolutional layers to extract hierarchical features, the Feature Pyramid Network (FPN) enhances detection at different scales.
-The network head performs final predictions, incorporating convolutional blocks and upsample blocks to refine features [30].
-III-B5 YOLOv9
-YOLOv9, developed by Chien-Yao Wang, I-Hau Yeh, and Hong-Yuan Mark Liao, uses the Information Bottleneck Principle and Reversible Functions to preserve essential data across the network’s depth, ensuring reliable gradient generation and improved model convergence and performance [70].
-Reversible functions, which can be inverted without loss of information, are another cornerstone of YOLOv9’s architecture.
-This property allows the network to retain a complete information flow, enabling more accurate updates to the model’s parameters.
-Moreover, YOLOv9 offers five scaled versions for different uses, focusing on lightweight models, which are often under-parameterized and prone to losing significant information during the feedforward process.
-Programmable Gradient Information (PGI) is a significant advancement introduced in YOLOv9.
-PGI is a method that dynamically adjusts the gradient information during training to optimize learning efficiency.
-By selectively focusing on the most informative gradients, PGI helps preserve crucial information that might be lost in lightweight models.
-This advancement ensures the model retains the essential features for accurate object detection, improving overall performance.
-In addition, YOLOv9 incorporates Gradient Enhanced Lightweight Architecture Network (GELAN), a new architectural advancement designed to improve parameter utilization and computational efficiency as illustrated in Figure 7.
-GELAN achieves this by optimizing the computational pathways within the network, allowing for better resource management and adaptability to various applications without compromising speed or accuracy.
+The network head performs final predictions, incorporating convolutional blocks and upsample blocks to refine features [35]
+3.2.5 YOLOv9
 Figure 8: YOLOv9 architecture featuring CSPNet, ELAN, and GELAN modules.
 CSPNet optimizes gradient flow and reduces computational complexity via feature map partitioning.
-ELAN enhances learning efficiency by linearly aggregating features, and GELAN extends this concept by integrating features from various depths and pathways [70].
-III-B6 YOLOv10
-YOLOv10, developed by Tsinghua University researchers, builds upon previous models’ strengths with key innovations [68].
-The architecture has an enhanced Cross Stage Partial Network (CSPNet) backbone for improved gradient flow and reduced computational redundancy, as seen in Figure 8.
-The network is structured into three main components: the backbone, the neck, and the detection head.
-The neck includes Path Aggregation Network (PAN) layers for effective multiscale feature fusion.
-PAN is designed to enhance information flow by aggregating features from different layers, enabling the network to better capture and combine details at various scales, which is crucial for detecting objects of different sizes.
-At the same time, the One-to-Many Head generates multiple predictions per object during training to provide rich supervisory signals and improve learning accuracy.
-Five scaled versions are available, from nano to extra-large.
-For inference, the One-to-One Head generates a single best prediction per object, eliminating the need for Non-Maximum Suppression (NMS).
-By removing the need for NMS, YOLOv10 reduces latency and improves the post-processing speed.
-In addition, YOLOv10 includes NMS-Free Training, which uses consistent dual assignments to reduce inference latency, and a model design that optimizes various components from both efficiency and accuracy perspectives.
-This includes lightweight classification heads, spatial-channel decoupled downsampling, and rank-guided block design.
-In addition, the model incorporates large-kernel convolutions and partial self-attention modules to enhance performance without significant computational costs.
+ELAN enhances learning efficiency by linearly aggregating features, and GELAN extends this concept by integrating features from various depths and pathways [82]
+YOLOv9, developed by Chien-Yao Wang, I-Hau Yeh, and Hong-Yuan Mark Liao, enhances object detection using the Information Bottleneck Principle and Reversible Functions, ensuring efficient gradient propagation and improved model convergence [82].
+Reversible functions enable lossless information retention, benefiting lightweight models prone to under-parameterization.
+A key advancement, Programmable Gradient Information (PGI), dynamically adjusts gradient flow during training, prioritizing informative gradients to prevent feature loss.
+Additionally, YOLOv9 integrates Gradient Enhanced Lightweight Architecture Network (GELAN), optimizing computational pathways for better parameter utilization, high-speed inference, and accurate detection, as illustrated in Figure 8.
+With five scalable versions, YOLOv9 balances efficiency and feature retention, making it highly adaptable for various real-world applications.
+3.2.6 YOLOv10
+YOLOv10, developed by Tsinghua University researchers, introduces key innovations, enhancing gradient flow and computational efficiency through an improved Cross Stage Partial Network (CSPNet) backbone [79].
+The architecture, illustrated in Figure 9, consists of three main components: the backbone, the neck, and the detection head.
+The Path Aggregation Network (PAN) neck improves multi-scale feature fusion, allowing for better detection of objects at different sizes.
+The One-to-Many Head generates multiple predictions per object during training, enhancing learning accuracy.
+For inference, YOLOv10 replaces Non-Maximum Suppression (NMS) with a One-to-One Head, producing a single best prediction per object, reducing latency and post-processing time.
+Additionally, NMS-Free Training leverages dual assignments, optimizing both speed and accuracy.
+Other enhancements include lightweight classification heads, spatial-channel decoupled downsampling, large-kernel convolutions, and partial self-attention modules, all improving efficiency without significantly increasing computational costs.
+YOLOv10 is available in five scaled versions, from nano to extra-large, catering to diverse use cases.
 Figure 9: YOLOv10 architecture showcasing the dual label assignment strategy for improving accuracy.
-The PAN layer enhances feature representation alongside one-to-many head for regression and classification tasks and one-to-one head for precise localization [68].
-III-B7 YOLO11
+The PAN layer enhances feature representation alongside one-to-many head for regression and classification tasks and one-to-one head for precise localization [79]
+3.2.7 YOLO11
+Figure 10: YOLO11 architecture showcasing the new C3k2 blocks and the C2PSA module [33] [54]
 YOLO11 [33] is one of the latest additions to the YOLO series developed by Ultralytics, building upon the developments of its predecessors, especially YOLOv8.
 This iteration offers five scaled models from nano to extra large, catering to various applications.
-Like YOLOv8, YOLO11 includes numerous applications such as object detection, instance segmentation, image classification, pose estimation, and oriented object detection (OBB).
+Like YOLOv8, YOLO11 includes numerous applications such as object detection, instance segmentation, image classification, pose estimation, and OBB.
 Key improvements in YOLO11 include the introduction of the Cross-Stage Partial with Self-Attention (C2PSA) module, as seen in Figure 10, which combines the benefits of cross-stage partial networks with self-attention mechanisms.
@@
 This block uses a smaller kernel, retaining accuracy while improving efficiency and speed.
-Figure 9: YOLO11 architecture showcasing the new C3k2 blocks and the C2PSA module. [28] [46].
-III-B8 YOLOv12
+3.2.8 YOLOv12
+Figure 11: YOLOv12 architecture showcasing the new Area Attention (A2) module and Residual Efficient Layer Aggregation Networks (R-ELAN)
 YOLOv12 [73] is the latest evolution in the YOLO series, introducing an attention-centric design that significantly enhances both speed and accuracy.
-It continues the trend of offering five scalable models (Nano to Extra Large), making it adaptable for a wide range of applications such as object detection, instance segmentation, and oriented object detection (OBB).
+It continues the trend of offering five scalable models (Nano to Extra Large), making it adaptable for a wide range of applications such as object detection, instance segmentation, and OBB.
 As seen in Figure 11, YOLOv12 introduces the Area Attention (A2) module, which maintains a large receptive field while drastically reducing computational complexity, allowing the model to enhance speed without compromising accuracy.
@@
 It replaces linear layers with convolutional layers and batch normalization, maximizing computational efficiency and achieving state-of-the-art latency-accuracy trade-offs.
-Figure 10: YOLOv12 architecture showcasing the new Area Attention (A2) module and Residual Efficient Layer Aggregation Networks (R-ELAN). [63]
-III-C Hardware and Software Setup
-Table IV presents the libraries and software versions used throughout this study.
-The experiments were conducted with Python 3.12, Ubuntu 22.04, CUDA 12.5, and cuDNN 8.9.7 for GPU acceleration.
+3.3 Hardware and Software Setup
+The experiments were conducted with Python 3.12, Ubuntu 22.04, CUDA 12.4, and cuDNN 8.9.7 for GPU acceleration.
 Ultralytics 8.2.55 was used for model training, while WandB 0.17.4 was utilized for tracking experiments.
-To ensure a fair comparison, similar hyperparameters were used across all models, as outlined in Table V.
-The key hyperparameters include 100 epochs, the AdamW optimizer, a batch size of 16, and an image size of (640, 640).
-The learning rate was initially set to 0.0001 and increased to 0.01, while the dropout rate was set to 0.15.
-The dataset was split into 70% for training, 20% for validation, and 10% for testing with no data augmentation techniques applied to any dataset.
+To ensure a fair comparison, similar hyperparameters were used across all models, as outlined in Table 3.
 The experiments were carried out using 2 NVIDIA RTX 4090 GPUs, each with 16,384 CUDA cores.
-TABLE IV: Software setup
-Name Version
-Python 3.12
-Ubuntu 22.04
-CUDA 12.5
-cuDNN 8.9.7
-Ultralytics 8.2.55
-WandB 0.17.4
-TABLE V: Table of parameters
+Table 3: Table of parameters
 Parameter Value
@@
 Data Split (70, 20, 10)
-III-D Metrics
-This study evaluates the performance of YOLO models using three primary metrics: accuracy, computational efficiency, and size.
-The accuracy metrics include Precision, Recall, Mean Average Precision at an Intersection over Union (IoU) threshold of 0.50 (mAP50), and Mean Average Precision at an Intersection over Union (IoU) threshold of 0.50-0.95 (mAP50-95).
-Precision [44] measures the ratio of correctly predicted observations to the total predicted observations, thus highlighting the occurrence of False Positives.
-Conversely, Recall [44] measures the ratio of correctly predicted observations to all actual observations, thus emphasizing the occurrence of False Negatives.
-Both mAP50 and mAP50-95 [75] provide a comprehensive summary of Precision and Recall.
-While mAP50 calculates the Mean Average Precision at an IoU threshold of 0.50, mAP50-95 extends this calculation across multiple IoU thresholds from 0.50 to 0.95, with 0.05 step size.
-Regarding computational efficiency metrics, Preprocessing Time, Inference Time, and Postprocessing Time will be utilized to evaluate the model’s speed.
-Preprocessing Time refers to the duration taken to prepare raw data for input into the model.
-Inference Time is the duration required for the model to process the input data and generate predictions.
-Postprocessing Time denotes the time needed to convert the model’s raw predictions into a final, usable format.
-These metrics were measured using a sample of images for testing after training the models.
-Additionally, the Giga Floating-Point Operations Per Second (GFLOPs) measure the computational power for the model training, reflecting its efficiency.
-In contrast, the size metric reflects the actual disk size of the model and the number of its parameters.
-These metrics are essential for providing a comprehensive overview of YOLO models’ performance, allowing for effective comparison and evaluation.
-These metrics allows a thorough assessment of the accuracy and efficiency of different YOLO models, ensuring a robust benchmark for their performance and applications in various real-world scenarios.
-IV Benchmark Results and Discussion
-IV-A Results
-IV-A1 Traffic Signs Dataset
-TABLE VI: Evaluation results for the Traffic Signs Dataset.
-Versions Precision Recall mAP50 mAP50-95
-Preprocess
-Time
-Inference
-Time
-Postprocess
-Time
-Total
-Time
-GFLOPs Size
-YOLOv3u 0.75 0.849 0.874 0.781 0.7 8.5 0.4 9.6 207.86 282.4
-YOLOV3u tiny 0.845 0.667 0.772 0.682 1.4 0.7 0.3 2.4 24.44 19
-YOLOv5un 0.805 0.679 0.749 0.665 0.6 6.6 0.4 7.6 5.65 7.1
-YOLOv5us 0.85 0.777 0.827 0.744 0.5 7.8 0.4 8.7 18.58 23.9
-YOLOv5um 0.849 0.701 0.83 0.744 1.1 9.5 0.4 11 50.54 64.1
-YOLOv5ul 0.831 0.836 0.886 0.799 0.6 9.7 0.4 10.7 106.85 134.9
-YOLOv5ux 0.863 0.795 0.867 0.777 1.1 9.8 0.4 11.3 195.2 246.3
-YOLOv8n 0.749 0.688 0.777 0.689 0.6 6.8 0.4 7.8 6.55 8.1
-YOLOv8s 0.766 0.788 0.806 0.718 0.6 7.8 0.4 8.8 22.59 28.6
-YOLOv8m 0.838 0.805 0.845 0.763 1.6 9.1 0.4 11.1 52.12 78.9
-YOLOv8l 0.771 0.789 0.853 0.767 0.6 9.2 0.4 10.2 87.77 165
-YOLOv8x 0.902 0.744 0.874 0.78 0.6 9.4 0.4 10.4 136.9 257.7
-YOLOv9t 0.792 0.748 0.812 0.731 0.5 10 0.4 10.9 4.93 7.7
-YOLOv9s 0.763 0.81 0.828 0.75 0.6 11.1 0.4 12.1 15.33 26.8
-YOLOv9m 0.864 0.796 0.864 0.784 1 12.1 0.4 13.5 40.98 76.7
-YOLOv9c 0.827 0.807 0.852 0.769 1.3 11.6 0.4 13.3 51.8 102.6
-YOLOv9e 0.819 0.824 0.854 0.764 0.8 16.1 0.4 17.3 117.5 189.4
-YOLOv10n 0.722 0.602 0.722 0.64 1 0.8 0.2 2 5.59 8.3
-YOLOv10s 0.823 0.742 0.834 0.744 1.2 1.1 0.2 2.5 15.9 24.7
-YOLOv10m 0.834 0.843 0.88 0.781 1.2 2.4 0.2 3.8 32.1 63.8
-YOLOv10b 0.836 0.764 0.859 0.765 1 3.1 0.2 4.3 39.7 98.4
-YOLOv10l 0.873 0.807 0.866 0.771 1.1 3.8 0.2 5.1 50 126.8
-YOLOv10x 0.773 0.854 0.88 0.787 1 6.3 0.2 7.5 61.4 170.4
-YOLO11n 0.768 0.695 0.757 0.668 1.2 0.6 0.4 2.2 5.35 6.4
-YOLO11s 0.819 0.758 0.838 0.742 1.2 1 0.4 2.6 18.4 21.4
-YOLO11m 0.898 0.826 0.893 0.795 1.2 2.4 0.4 4 38.8 67.9
-YOLO11l 0.862 0.839 0.889 0.794 1.2 3 0.4 4.6 49 86.8
-YOLO11x 0.819 0.816 0.885 0.784 0.9 6.1 0.4 7.4 109 194.8
-YOLOv12n 0.883 0.684 0.802 0.708 1.2 4.6 0.4 6.2 5.34 6.4
-YOLOv12s 0.761 0.839 0.877 0.78 1.6 4.3 0.5 6.4 18.1 21.4
-YOLOv12m 0.868 0.811 0.89 0.79 0.9 4.9 0.4 6.2 39 67.3
-YOLOv12l 0.826 0.836 0.877 0.781 0.9 5.5 0.4 6.8 51.2 89.6
-YOLOv12x 0.832 0.849 0.889 0.793 1 9.5 0.5 11 114 198.9
-Table VI presents a comparative analysis of the YOLO algorithms’ performance on the Traffic Signs dataset, evaluated based on accuracy, computational efficiency, and model size.
+3.4 Metrics
+This study evaluates YOLO models based on accuracy, computational efficiency, and size.
+Accuracy metrics include Precision, Recall, mAP50, and mAP50-95.
+Precision measures the ratio of correctly predicted observations to total predictions, indicating False Positives, while Recall highlights False Negatives by comparing correct predictions to actual observations [52].
+mAP50 calculates Mean Average Precision at an IoU threshold of 0.50, whereas mAP50-95 extends this across thresholds from 0.50 to 0.95 with a 0.05 step size [88].
+Computational efficiency is assessed using Preprocessing Time, Inference Time, and Postprocessing Time.
+Preprocessing Time refers to data preparation, Inference Time measures prediction generation, and Postprocessing Time converts raw outputs into final results.
+Additionally, GFLOPs indicate computational power, while model size reflects storage requirements and parameter count.
+These metrics enable a comprehensive performance comparison, ensuring a robust benchmark for real-world applications.
+4 Benchmark Results
+4.1 Results
+4.1.1 Traffic Signs Dataset
+Figure 12: Evaluation results for the traffic signs dataset.
+Figure 12 presents a comparative analysis of the YOLO algorithms’ performance on the Traffic Signs dataset, evaluated based on accuracy, computational efficiency, and model size.
 The Traffic Signs dataset is a medium-sized dataset with varied object sizes, making it favorable for benchmarking.
-The results highlight the effectiveness of YOLO models in detecting traffic signs, demonstrating a range of precision.
-The highest mAP50-95 was 0.799, while the lowest recorded precision was 0.64.
-On the other hand, the highest mAP50 is 0.893 while the lowest is 0.722.
-The substantial gap between the mAP50 and mAP50-95 results suggests that the models have difficulties in uniformly handling traffic signs with different sizes at higher thresholds, reflecting areas for potential improvement in their detection algorithms.
 Accuracy:
@@
 This is followed by YOLO11m with a mAP50-95 of 0.795 and YOLO11l with a mAP50-95 of 0.794.
-In contrast, YOLOv10n exhibits the lowest precision, with a mAP50 of 0.722 and a mAP50-95 of 0.64, closely followed by YOLOv5un with a mAP50-95 of 0.665, as evidenced by the data points in Figure 11.
-Figure 11: mAP50 and mAP50-95 YOLO results on traffic signs dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
-Figure 12: Precision vs. Recall based on size on traffic signs dataset, with larger circles indicating larger model sizes.
-Precision and Recall:
-Figure 12 elucidates the trade-off between precision and recall taking the size of the models into consideration.
-Models such as YOLO11m, YOLO10l, YOLOv12m, YOLOv9m, YOLOv5ux, and YOLO11l exhibit high precision and recall, specifically with YOLO11m achieving a precision of 0.898 and a recall of 0.826 while having a size of 67.9Mb, and YOLOv12m achieving a precision of 0.868 and a recall of 0.811 with a significantly smaller size (39 Mb).
-In contrast, smaller models such as YOLOv10n (precision 0.722, recall 0.602), YOLOv8n (precision 0.749, recall 0.688), and YOLO11n (precision 0.768, recall 0.695) underperform in both metrics.
-This underscores the superior performance of larger models on the Traffic Signs dataset.
-Moreover, the high precision (0.883) and low recall (0.684) of YOLOv12n indicate a propensity for false negatives, while YOLOv3u’s high recall (0.849) and low precision (0.75) suggest a tendency for false positives.
-Figure 13: Total processing time and GFLOPs count results on traffic signs dataset.
+In contrast, YOLOv10n exhibits the lowest precision, with a mAP50 of 0.722 and a mAP50-95 of 0.64, closely followed by YOLOv5un with a mAP50-95 of 0.665.
+Figure 13: mAP50 and mAP50-95 YOLO results on traffic signs dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score
+Figure 14: Precision vs. Recall based on size on traffic signs dataset, with larger circles indicating larger model sizes
+Figure 15: Total processing time and GFLOPs count results on traffic signs dataset
 Computational Efficiency:
-In terms of computational efficiency, YOLOv10n is the most efficient, with a processing time of 2ms per image and a GFLOPs count of 8.3, as shown in Figure 13.
-YOLO11n closely trails this at 2.2ms with a 6.4 GFLOPs count, and YOLOv3u-tiny with a processing time of 2.4ms and a GFLOPs count of 19, making it relatively computationally inefficient compared to the other fast models.
+Regarding computational efficiency, YOLOv10n is the most efficient, with a processing time of 2ms per image and a GFLOPs count of 8.3, as shown in Figure 15.
+YOLO11n closely trails this at 2.2ms with a 6.4 GFLOPs count, and YOLOv3u tiny with a processing time of 2.4ms and a GFLOPs count of 19, making it relatively computationally inefficient compared to the other fast models.
 However, the data indicates that YOLOv9e, YOLOv9m, YOLOv9c, and YOLOv9s are the least efficient, with inference times of 16.1ms, 12.1ms, 11.6ms, and 11.1ms, and GFLOPs count of 189.4, 76.7, 102.6, and 26.8, respectively.
@@
 Overall Performance:
-When evaluating overall performance, which includes accuracy, size, and model efficiency, YOLO11m emerges as a consistently top-performing model.
-It achieves a mAP50-95 of 0.795, an inference time of 2.4ms, a model size of 38.8Mb, and a 67.9 GFLOPs count, as detailed in Figures 11, 13, and Table VI.
-This is followed by YOLO11l (mAP50-95 of 0.794, inference time of 4.6ms, size of 49Mb, and 86.8 GFLOPs count), and YOLOv10m (mAP50-95 of 0.781, inference time of 2.4ms, size of 32.1Mb, 63.8 GFLOPs count).
-These results highlight the robustness of these models in detecting traffic signs of various sizes while maintaining short inference times and small model sizes.
-On the contrary, YOLOv9e demonstrates poor performance overall, achieving a relatively low mAP50-95 of 0.764, an inference time of 16.1ms, a large model size of 117.5MB, and one of the highest GFLOPs counts at 198.4.
+When evaluating overall performance, which includes accuracy, size, and model efficiency, YOLO11m emerges as a consistently top-performing model as detailed in Figures 12, 13, 14, and 15.
+This is followed by YOLO11l and YOLOv10m showcasing the robustness of these models.
+On the contrary, YOLOv9e demonstrates poor performance overall in terms of accuracy, speed, GFLOPs, and model size.
 Additionally, YOLOv5um, YOLOv8m, and YOLOv8s performed suboptimally, further indicating the inconsistency in performance among earlier model families.
-Notably, the YOLO11 and YOLOv12 families significantly outperform other YOLO families in terms of accuracy and computational efficiency.
-Their models consistently surpass counterparts from the YOLOv3u, YOLOv5u, YOLOv8, YOLOv9, YOLOv10 families, demonstrating their ability to balance precision and speed effectively in the detection of objects with varying sizes.
-IV-A2 Africa Wildlife Dataset
-TABLE VII: Evaluation results for the Africa Wildlife dataset.
-Versions Precision Recall mAP50 mAP50-95
-Preprocess
-Time
-Inference
-Time
-Postprocess
-Time
-Total
-Time
-GFLOPs Size
-YOLOv3u 0.75 0.849 0.874 0.781 0.7 8.5 0.4 9.6 207.86 282.4
-YOLOV3u tiny 0.845 0.667 0.772 0.682 1.4 0.7 0.3 2.4 24.44 19
-YOLOv5un 0.805 0.679 0.749 0.665 0.6 6.6 0.4 7.6 5.65 7.1
-YOLOv5us 0.85 0.777 0.827 0.744 0.5 7.8 0.4 8.7 18.58 23.9
-YOLOv5um 0.849 0.701 0.83 0.744 1.1 9.5 0.4 11 50.54 64.1
-YOLOv5ul 0.831 0.836 0.886 0.799 0.6 9.7 0.4 10.7 106.85 134.9
-YOLOv5ux 0.863 0.795 0.867 0.777 1.1 9.8 0.4 11.3 195.2 246.3
-YOLOv8n 0.749 0.688 0.777 0.689 0.6 6.8 0.4 7.8 6.55 8.1
-YOLOv8s 0.766 0.788 0.806 0.718 0.6 7.8 0.4 8.8 22.59 28.6
-YOLOv8m 0.838 0.805 0.845 0.763 1.6 9.1 0.4 11.1 52.12 78.9
-YOLOv8l 0.771 0.789 0.853 0.767 0.6 9.2 0.4 10.2 87.77 165
-YOLOv8x 0.902 0.744 0.874 0.78 0.6 9.4 0.4 10.4 136.9 257.7
-YOLOv9t 0.792 0.748 0.812 0.731 0.5 10 0.4 10.9 4.93 7.7
-YOLOv9s 0.763 0.81 0.828 0.75 0.6 11.1 0.4 12.1 15.33 26.8
-YOLOv9m 0.864 0.796 0.864 0.784 1 12.1 0.4 13.5 40.98 76.7
-YOLOv9c 0.827 0.807 0.852 0.769 1.3 11.6 0.4 13.3 51.8 102.6
-YOLOv9e 0.819 0.824 0.854 0.764 0.8 16.1 0.4 17.3 117.5 189.4
-YOLOv10n 0.722 0.602 0.722 0.64 1 0.8 0.2 2 5.59 8.3
-YOLOv10s 0.823 0.742 0.834 0.744 1.2 1.1 0.2 2.5 15.9 24.7
-YOLOv10m 0.834 0.843 0.88 0.781 1.2 2.4 0.2 3.8 32.1 63.8
-YOLOv10b 0.836 0.764 0.859 0.765 1 3.1 0.2 4.3 39.7 98.4
-YOLOv10l 0.873 0.807 0.866 0.771 1.1 3.8 0.2 5.1 50 126.8
-YOLOv10x 0.773 0.854 0.88 0.787 1 6.3 0.2 7.5 61.4 170.4
-YOLO11n 0.768 0.695 0.757 0.668 1.2 0.6 0.4 2.2 5.35 6.4
-YOLO11s 0.819 0.758 0.838 0.742 1.2 1 0.4 2.6 18.4 21.4
-YOLO11m 0.898 0.826 0.893 0.795 1.2 2.4 0.4 4 38.8 67.9
-YOLO11l 0.862 0.839 0.889 0.794 1.2 3 0.4 4.6 49 86.8
-YOLO11x 0.819 0.816 0.885 0.784 0.9 6.1 0.4 7.4 109 194.8
-YOLOv12n 0.883 0.684 0.802 0.708 1.2 4.6 0.4 6.2 5.34 6.4
-YOLOv12s 0.761 0.839 0.877 0.78 1.6 4.3 0.5 6.4 18.1 21.4
-YOLOv12m 0.868 0.811 0.89 0.79 0.9 4.9 0.4 6.2 39 67.3
-YOLOv12l 0.826 0.836 0.877 0.781 0.9 5.5 0.4 6.8 51.2 89.6
-YOLOv12x 0.832 0.849 0.889 0.793 1 9.5 0.5 11 114 198.9
-The results in Table VII showcase the performance of the YOLO models on the Africa Wildlife dataset.
-This dataset contains large object sizes focusing on the ability of YOLO models to predict large objects and their risk of overfitting due to the size of the dataset.
-The models demonstrate robust accuracy across the board, with the highest-performing models achieving a mAP50-95 ranging from 0.832 to 0.725.
-This relatively shorter range reflects the effectiveness of the models in detecting and classifying large wildlife objects by maintaining high accuracy.
-Figure 14: mAP50 and mAP50-95 YOLO results on Africa wildlife dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
-Figure 15: Precision vs. Recall based on size results on Africa wildlife dataset, with larger circles indicating larger model sizes.
-Figure 16: Total processing time and GFLOPs count results on Africa wildlife dataset.
+Notably, the YOLO11 and YOLOv12 families significantly outperform other YOLO families regarding accuracy and computational efficiency.
+Their models consistently surpass counterparts from the YOLOv3u, YOLOv5u, YOLOv8, YOLOv9, and YOLOv10 families, demonstrating their ability to balance precision and speed effectively in the detection of objects with varying sizes.
+4.1.2 Africa Wildlife Dataset
+Figure 16: Evaluation results for the Africa wildlife dataset.
+The results in Figure 16 showcase the performance of the YOLO models on the Africa Wildlife dataset.
+This dataset contains large object sizes, focusing on the ability of YOLO models to predict large objects and their risk of overfitting due to the size of the dataset.
+Figure 17: mAP50 and mAP50-95 YOLO results on Africa wildlife dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score
+Figure 18: Precision vs. Recall based on size results on Africa wildlife dataset, with larger circles indicating larger model sizes
+Figure 19: Total processing time and GFLOPs count results on Africa wildlife dataset
 Accuracy:
@@
 These results highlight the YOLOv9 family’s ability to effectively learn patterns from a small sample of images, making it particularly suited for smaller datasets.
-In contrast, YOLOv5un, YOLOv10n, and YOLOv3u-tiny show lower mAP50-95 scores of 0.791, 0.786, and 0.725, indicating their limitations in accuracy.
+In contrast, YOLOv5un, YOLOv10n, and YOLOv3u tiny show lower mAP50-95 scores of 0.791, 0.786, and 0.725, indicating their limitations in accuracy.
 The underperformance of larger models like YOLO11x, YOLOv5ux, YOLOv5ul, and YOLOv10l can be attributed to overfitting, especially given the small dataset size.
-Precision and Recall:
-Figure 15 reveals that YOLO8l and YOLO11l achieve the highest precision and recall, with values of 0.942 and 0.937 for precision, and 0.898 and 0.896 for recall, respectively.
-Notably, YOLOv8n achieves similar results (0.932 for precision, 0.908 for recall) with a compact size of 6.55Mb, demonstrating its efficiency.
-In contrast, YOLOv3u and YOLOv5ul exhibit lower precision and recall scores (0.91 and 0.88 for YOLOv3u, 0.916 and 0.881 for YOLOv5ul), despite their larger sizes (204.86Mb for YOLOv3u, 106.85Mb for YOLOv5ul), which may be attributed to overfitting issues.
 Computational Efficiency:
-YOLOv10n, YOLOv8n, and YOLOv3u-tiny are the fastest models, achieving processing times of 2ms and 1.8ms, with GFLOPs counts of 8.2 and 19.1, respectively.
+YOLOv10n, YOLOv8n, and YOLOv3u tiny are the fastest models, achieving processing times of 2ms and 1.8ms, with GFLOPs counts of 8.2 and 19.1, respectively.
 The first two models share the same processing speed and GFLOPs count, as showcased in Figure 19.
@@
 Overall Performance:
-YOLOv9t and YOLOv9s consistently excel across all metrics, delivering high accuracy while maintaining small model sizes, low GFLOPs, and short inference times, as shown in Table VII, and Figures 15 and 16.
+YOLOv9t and YOLOv9s consistently excel across all metrics, delivering high accuracy while maintaining small model sizes, low GFLOPs, and short inference times, as shown in Figures 16, 18, 18, and 19.
 This demonstrates the robustness of YOLOv9’s smaller models and their effectiveness on small datasets.
@@
 Most large models underperformed on this dataset, with the exception of YOLOv10x, which benefited from a modern architecture that prevents overfitting.
-TABLE VIII: Evaluation results for the ships and vessels dataset.
-Versions Precision Recall mAP50 mAP50-95
-Preprocess
-Time
-Inference
-Time
-Postprocess
-Time
-Total
-Time
-Size GFLOPs
-YOLOv3u 0.679 0.534 0.519 0.298 0.8 6.2 0.3 7.3 207.86 282.5
-YOLOV3u tiny 0.647 0.511 0.489 0.273 1 0.7 0.3 2 24.44 18.9
-YOLOv5un 0.635 0.532 0.514 0.298 1.5 0.6 0.3 2.4 5.65 7.2
-YOLOv5us 0.653 0.541 0.518 0.299 1.2 0.8 0.3 2.3 18.58 24
-YOLOv5um 0.667 0.541 0.526 0.308 0.9 2.1 0.3 3.3 50.54 64
-YOLOv5ul 0.654 0.545 0.525 0.305 0.9 3.3 0.3 4.5 106.85 134.8
-YOLOv5ux 0.668 0.555 0.531 0.309 0.8 6.7 0.3 7.8 195.2 246.2
-YOLOv8n 0.655 0.533 0.515 0.297 1.5 0.5 0.3 2.3 6.55 8
-YOLOv8s 0.647 0.545 0.518 0.301 1.1 1 0.3 2.4 22.59 28.5
-YOLOv8m 0.669 0.547 0.525 0.302 0.8 2.5 0.3 3.6 52.12 79
-YOLOv8l 0.659 0.551 0.526 0.303 0.9 3.9 0.3 5.1 87.77 165
-YOLOv8x 0.655 0.55 0.529 0.306 0.8 7.1 0.3 8.2 136.9 257.7
-YOLOv9t 0.647 0.516 0.512 0.3 1.4 1.1 0.3 2.8 4.93 7.5
-YOLOv9s 0.655 0.552 0.522 0.308 1.4 1.2 0.3 2.9 15.33 26.9
-YOLOv9m 0.668 0.551 0.529 0.307 1.1 2.7 0.3 4.1 40.98 76.8
-YOLOv9c 0.663 0.547 0.523 0.303 1.2 3.4 0.3 4.9 51.8 102.4
-YOLOv9e 0.667 0.537 0.524 0.308 1.1 7.6 0.3 9 117.5 189.5
-YOLOv10n 0.584 0.487 0.506 0.31 1.4 0.8 0.2 2.4 5.59 8.2
-YOLOv10s 0.586 0.511 0.515 0.319 1.1 1.1 0.2 2.4 15.9 24.4
-YOLOv10m 0.588 0.517 0.522 0.322 1 2.4 0.1 3.5 32.1 63.4
-YOLOv10b 0.603 0.509 0.523 0.319 1.1 3.2 0.1 4.4 39.7 97.9
-YOLOv10l 0.601 0.511 0.522 0.322 1.1 3.8 0.1 5 50 126.3
-YOLOv10x 0.6 0.523 0.526 0.321 1 6.3 0.2 7.5 61.4 169.8
-YOLO11n 0.574 0.51 0.505 0.311 1.5 0.7 0.3 2.5 5.35 6.3
-YOLO11s 0.585 0.535 0.521 0.323 1.3 1 0.3 2.6 18.4 21.3
-YOLO11m 0.588 0.541 0.53 0.325 1 2.4 0.3 3.7 38.8 67.6
-YOLO11l 0.596 0.531 0.528 0.325 1.1 3 0.4 4.5 49 86.6
-YOLO11x 0.596 0.538 0.529 0.327 0.8 6.1 0.3 7.2 109 194.4
-YOLOv12n 0.571 0.505 0.505 0.307 1.1 4.7 0.4 6.2 5.34 6.3
-YOLOv12s 0.58 0.529 0.516 0.319 1.6 5 0.4 7 18.1 21.2
-YOLOv12m 0.584 0.543 0.525 0.325 1.1 4.6 0.4 6.1 39 67.1
-YOLOv12l 0.588 0.532 0.522 0.323 2.6 5.2 0.4 8.2 51.2 88.5
-YOLOv12x 0.586 0.538 0.52 0.321 1.2 9.3 0.4 10.9 114 198.5
-Figure 17: mAP50 and mAP50-95 YOLO results on ships and vessel dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score.
-Figure 18: Precision vs. Recall based on size results on ships and vessels dataset, with larger circles indicating larger model sizes.
-Figure 19: Total processing time and GFLOPs count results on ships and vessels dataset.
-IV-A3 Ships and Vessels Dataset
-Table VIII presents the performance of YOLO models on the Ships and Vessels dataset, a large dataset featuring tiny objects with varying orientations.
-Overall, the models demonstrated moderate effectiveness in detecting ships and vessels, with mAP50-95 ranging from 0.273 to 0.327.
-This performance suggests that YOLO algorithms may face challenges in accurately detecting smaller objects with different orientations.
+Figure 20: Evaluation results for the ships and vessels dataset
+Figure 21: mAP50 and mAP50-95 YOLO results on ships and vessel dataset. Each model is represented by two bars: the left bar shows the mAP50 score, while the right bar represents the mAP50-95 score
+Figure 22: Precision vs. Recall based on size results on ships and vessels dataset, with larger circles indicating larger model sizes
+Figure 23: Total processing time and GFLOPs count results on ships and vessels dataset
+4.1.3 Ships and Vessels Dataset
+Figure 20 presents the performance of YOLO models on the Ships and Vessels dataset, a large dataset featuring tiny objects with varying orientations.
 Accuracy:
@@
 Among the models, YOLO11x achieved the highest accuracy, with a mAP50 of 0.529 and a mAP50-95 of 0.327, closely followed by YOLO11l, YOLO11m, and YOLO11s, which recorded mAP50 values of 0.529, 0.528, and 0.53, and mAP50-95 values of 0.327, 0.325, and 0.325, respectively.
-These results highlight the robustness of the YOLO11 family in detecting small and tiny objects.
 In contrast, YOLOv3u-tiny, YOLOv8n, YOLOv3u, and YOLOv5n exhibited the lowest accuracy, with mAP50 scores of 0.489, 0.515, 0.519, and 0.514, and mAP50-95 scores of 0.273, 0.297, 0.298, and 0.298, respectively.
-This suggests the outdated YOLOv3u architecture and the potential underfitting of smaller models with large size of datasets.
-Precision and Recall:
-Figure 18 indicates that YOLOv5ux outperformed other models, achieving a precision of 0.668 and a recall of 0.555.
-It was closely followed by YOLOv9m (precision of 0.668, recall of 0.551) and YOLOv8m (precision of 0.669, recall of 0.525), both of which are significantly smaller in size (40.98 Mb for YOLOv9m and 52.12 Mb for YOLOv8m).
-In contrast, YOLOv12n and YOLO11n exhibited lower performance, with precisions of 0.571 and 0.574 and recalls of 0.505 and 0.51, respectively, likely due to underfitting issues.
-Generally, YOLOv12 and YOLO11 models tended to produce false positives, reflected in their low precision and high recall.
-Meanwhile, YOLOv10 underperformed in both precision and recall, despite being one of the newest models in the YOLO family.
+This suggests the outdated YOLOv3u architecture and the potential underfitting of smaller models with large datasets.
 Computational Efficiency:
-As illustrated in Figure 19, YOLOv3u-tiny achieved the fastest processing time at 2 ms, closely followed by YOLOv8n and YOLOv5un, both recording 2.3 ms.
+As illustrated in Figure 23, YOLOv3u tiny achieved the fastest processing time at 2 ms, closely followed by YOLOv8n and YOLOv5un, recording 2.3 ms.
 YOLOv10 and YOLO11 models also excelled in speed, with YOLOv10n and YOLO11n achieving rapid inference times of 2.4 ms and 2.5 ms, along with GFLOPs counts of 8.2 and 6.3, respectively.
@@
 Overall Performance:
-The results in Table VIII and Figures 17, 18, and 19 demonstrate that YOLO11s and YOLOv10s excelled in accuracy while maintaining compact sizes, low GFLOPs, and quick processing times.
+The results in Figures 20, 21, 22, and 23 demonstrate that YOLO11s and YOLOv10s excelled in accuracy while maintaining compact sizes, low GFLOPs, and quick processing times.
 In contrast, YOLOv3u, YOLOv8x, and YOLOv8l fell short of expectations despite their larger sizes and longer processing times.
 These findings highlight the robustness and reliability of the YOLO11 family, particularly in improving the YOLO family’s performance in detecting small and tiny objects while ensuring efficient processing.
-Additionally, the results reveal the underperformance of YOLOv9 models when faced with large datasets and small objects, despite their modern architecture.
-IV-B Discussion
-Based on the performance of the models across the three datasets, we ranked them by accuracy, speed, GFLOps count, and size, as shown in Table IX to facilitate a comprehensive evaluation.
-For accuracy, the mAP50-95 metric was employed due to its capacity to assess models across a range of IoU thresholds, thus providing a detailed insight into each model’s performance.
-For speed, models were sorted based on the total processing time, which encompasses preprocessing, inference, and postprocessing durations.
-The rankings range from Rank 1, indicating the highest performance, to Rank 28, denoting the lowest, with the respective rankings highlighted in bold within the table.
-TABLE IX: Overall ranking of YOLO algorithms
+Additionally, the results reveal the underperformance of YOLOv9 models when faced with large datasets and small objects despite their modern architecture.
+5 Discussion
+Models were ranked by accuracy, speed, GFLOPs count, and size across three datasets, as shown in Figure 24.
+To ensure fair comparisons, models were categorized by scale.
+YOLOv3u tiny was classified as small, YOLOv9t as nano, YOLOv9c as medium, YOLOv9c and YOLOv10b as large, and YOLOv9e as extra-large.
+Additionally, family-wide performance was analyzed by averaging rankings across all scales to compare overall performance as seen in Table 4.
+Accuracy was measured using mAP50-95 for a comprehensive performance assessment, while speed rankings were based on total processing time .
+Rankings range from Rank 1 (best) to the lowest performer, with results highlighted in bold.
+This approach ensures a fair evaluation of architectural improvements and computational trade-offs across model sizes and families.
+Figure 24: Overall ranking of YOLO algorithms based on size
+Table 4: Overall ranking of YOLO algorithms based on family
 Version Acc. Speed GFLOPs Size
-YOLOv3u tiny 33 1 7 33
-YOLOv3u 25 28 33 13
-YOLOv5un 32 6 3 5
-YOLOv5us 29 8 10 11
-YOLOv5um 21 15 15 21
-YOLOv5ul 18 21 25 27
-YOLOv5ux 22 31 31 32
-YOLOv8n 31 5 5 6
-YOLOv8s 26 9 13 12
-YOLOv8m 19 18 19 24
-YOLOv8l 17 23 26 26
-YOLOv8x 11 30 32 31
-YOLOv9t 23 14 4 1
-YOLOv9s 10 16 12 7
-YOLOv9m 5 22 18 18
-YOLOv9c 13 26 23 23
-YOLOv9e 16 33 28 30
-YOLOv10n 30 2 6 4
-YOLOv10s 23 3 11 8
-YOLOv10m 9 9 14 14
-YOLOv10b 11 12 22 17
-YOLOv10l 14 16 24 20
-YOLOv10x 2 26 27 25
-YOLO11n 27 3 1 3
-YOLO11s 19 6 8 10
-YOLO11m 1 11 17 15
-YOLO11l 2 13 20 19
-YOLO11x 8 23 29 28
-YOLOv12n 27 19 1 2
-YOLOv12s 7 25 8 9
-YOLOv12m 5 20 16 16
-YOLOv12l 4 28 21 22
-YOLOv12x 15 32 30 29
-The analysis of Table IX yields critical observations:
-IV-B1 Accuracy
-YOLO11m consistently emerged as a top performer, frequently ranking among the highest, closely followed by YOLOv10x, YOLO11l, YOLOv9m, and YOLO11x.
-This underscores the robust performance of the YOLO11 family across varying IoU thresholds and object sizes, which can be attributed to their use of C2PSA for the preservation of contextual information, leading to improved convergence and overall performance.
-The inclusion of two small kernel convolutions and partial self-attention modules in these models helped capture more spatial features, leading to better accuracy in detecting both large and small objects.
-Additionally, the fine-tuning of anchor boxes and multi-scale training has enhanced the detection of objects at different resolutions, which is crucial in achieving high accuracy across diverse datasets.
-Conversely, YOLOv3u-tiny exhibited the lowest accuracy, particularly in the Africa Wildlife and Ships and Vessels datasets.
-YOLOv5un and YOLOv8n show slightly better but still sub-par results.
-This suggests that YOLO11 models are currently the most reliable for applications demanding high accuracy.
-The YOLOv3u-tiny model’s underperformance can be attributed to its older architecture, which lacks the optimizations found in newer YOLO models.
-YOLOv3u is based on the Darknet-53 framework, which, although efficient for its time, struggles to capture complex patterns and fine details required for high-accuracy detection.
-Its relatively shallow architecture and absence of more recent enhancements such as attention mechanisms or advanced feature fusion techniques make it less effective in handling intricate object detection tasks, especially in datasets with a wide variety of object sizes, like those in the Africa Wildlife and Ships and Vessels datasets.
-In contrast, the YOLOv9 models show effectiveness in detecting objects across various sizes and IoU thresholds, but they still struggle with small objects, particularly in the Ships and Vessels dataset.
-The YOLOv9 family benefits from improved training strategies, such as anchor box refinement and better scale handling, but may still fall short in detecting small objects in crowded or cluttered scenes due to its reliance on larger-scale feature maps and a less efficient processing pipeline for fine-grained features.
-The YOLOv10 family, despite its later introduction, exhibited relatively lower accuracy in the Traffic Signs and Africa Animals datasets, resulting in an average accuracy drop of 2.075% compared to the YOLOv9 models in those datasets.
-YOLOv10’s adoption of the One-to-One Head approach instead of Non-Maximum Suppression (NMS) for defining bounding boxes contributed to this decrease.
-The One-to-One Head approach relies on predicting a single bounding box per object, which can be a limitation in scenarios with overlapping or closely packed objects.
-NMS, on the other hand, is more effective at filtering out redundant predictions, which can improve detection accuracy in these types of datasets.
-Therefore, the One-to-One Head’s inability to handle multiple overlapping predictions adequately explains YOLOv10’s underperformance, especially in datasets like Traffic Signs, where multiple signs are often close together.
-Although the YOLOv12 family includes strong models like YOLOv12l, which ranked 4th in accuracy, it falls short compared to previous iterations, particularly considering its status as the newest YOLO version.
-It shows a -2.72% decline compared to YOLOv9 in the African Animals dataset and a -1% decrease in accuracy compared to YOLO11 in the Ships and Vessels dataset.
-This decline in accuracy is largely due to architectural choices.
-R-ELAN and area attention, while innovative, compromised global dependency modeling, impacting accuracy.
-Additionally, the adjusted MLP ratio prioritized certain computational aspects over feature representation depth, limiting its ability to capture intricate patterns.
-Consequently, YOLOv12 underperformed compared to the more balanced YOLOv11 and YOLOv9 families, demonstrating that newer architectural changes did not yield the expected gains in accuracy.
-Similarly, YOLOv8n, though it performed better than YOLOv3u, still showed lower accuracy than the top-performing models in specific datasets.
-This model demonstrated a good balance between speed and accuracy but struggled with intricate object recognition tasks.
-The simplicity of its architecture, without more advanced modules like self-attention or multi-scale training, likely held it back from achieving the high accuracy observed in models like YOLO11 and YOLOv9.
-Lastly, the YOLO11 models stand out for their ability to consistently achieve the highest accuracy.
-The use of advanced mechanisms like C2PSA and two kernel convolutions allowed for improved feature extraction across multiple scales, enhancing their ability to detect objects of varying sizes and complexities.
-These models’ robust performance across a variety of datasets, including challenging ones like Africa Wildlife and Ships and Vessels, demonstrates their ability to handle a wide range of object detection tasks with exceptional accuracy.
-IV-B2 Computational Efficiency
-YOLOv10n consistently outperformed other models in terms of speed and GFLOPs count, ranking among the top for speed across all datasets and 5th in terms of GFLOPs.
-Its use of CSPNet as backbone design and optimization strategies allowed for high processing speed without significantly compromising performance.
-YOLOv3u-tiny, YOLOv10s, and YOLO11n also demonstrated excellent computational efficiency, with their lightweight architectures enabling faster inference times and lower resource usage.
-Despite its accuracy advantages, YOLOv12 ranked as the worst family in terms of processing time and second in GFLOPs count.
-This is largely due to its attention-centric architecture, which enhances accuracy but significantly increases computational complexity.
-The integration of Area Attention and R-ELAN, while effective in capturing spatial features, introduces substantial overhead in feature extraction and aggregation, leading to slower inference times.
-Additionally, the reliance on FlashAttention contributes to inefficient memory access patterns, further impacting processing speed.
-Although these architectural choices improve contextual feature extraction, they hinder computational efficiency, making YOLOv12 less suitable for real-time applications compared to its predecessors.
-This performance trade-off illustrates that while attention mechanisms can elevate accuracy, they come at the cost of increased latency and computational demands.
-YOLOv5un, YOLOv5us, and YOLOv8n exhibited strong efficiency, providing a good balance between speed and accuracy.
-These models performed well in both inference times and GFLOPs, with YOLOv5un showing slightly better speed than YOLOv8n potentially due to the implementation of C2k (CSP with larger kernel sizes) in YOLOv8 architecture that increases accuracy but with a trade-off in computational efficiency.
-YOLOv8n, despite its good speed performance, had a larger size and computational cost compared to some other lightweight models.
-While YOLOv3u-tiny ranked highly in speed, its performance in accuracy was subpar, particularly in datasets like Africa Wildlife and Ships and Vessels.
-This highlights a critical trade-off: while speed is crucial for real-time applications, it becomes irrelevant if it severely compromises accuracy.
-YOLOv3u-tiny’s lightweight architecture, while optimized for faster inference, failed to detect objects accurately across varied and complex datasets, making it unsuitable for applications where precision is key.
-On the other hand, YOLOv3u struggled in both speed and accuracy, ranking poorly in both categories.
-Its reliance on the outdated Darknet-53 backbone contributed to this underperformance, as it was unable to meet modern detection challenges.
-The model’s low ranking in both computational efficiency and accuracy makes it less viable compared to newer architectures, such as YOLOv11 and YOLOv10, which offer better performance across both metrics.
-In contrast, YOLO11 models demonstrated significant speed improvements, thanks to the C3k2 block, which reduces computational overhead while maintaining high accuracy.
-These models were faster than YOLOv10 by 1.41% and YOLOv9 by 31% on average, making them ideal for real-time applications requiring rapid processing without sacrificing accuracy.
-While YOLOv9 models excelled in accuracy, their slower inference times made them less ideal for time-sensitive tasks.
-YOLOv10, slightly slower than YOLO11, offered a good balance between speed and accuracy, making it well-suited for real-time applications.
-Lastly, YOLOv5 and YOLOv8 families also performed well, with YOLOv5un ranking near the top for speed and efficiency, making it a solid choice for applications where both quick processing and solid accuracy are required.
-IV-B3 Model Size
-YOLOv9t emerged as the smallest model, ranking first across all three datasets in terms of model size thanks to its use of GELAN architecture leading to more efficient use of parameters.
-It is followed closely by YOLOv12n, YOLO11n, and YOLOv10n.
-Specifically, YOLOv10n benefits from the Spatial-Channel Decoupled Downsampling technique, which effectively reduces the model size while maintaining high accuracy, making it a highly efficient choice for resource-constrained environments.
-This technique allows for better utilization of parameters, enabling smaller models to perform similarly to larger counterparts.
-In contrast, YOLOv3u was the largest model, ranking poorly in terms of both size and accuracy.
-This inefficiency stems from its reliance on the outdated Darknet-53 architecture, which is larger and more computationally expensive compared to more modern designs.
-The size of YOLOv3u further exacerbates its subpar performance, as the model requires more resources and longer inference times, without corresponding improvements in detection accuracy.
-Its larger size makes it less suitable for deployment in environments where computational resources are limited or where real-time processing is essential.
-This comparison highlights the trend in newer YOLO versions, such as YOLOv9, YOLO10, and YOLO11, where optimizations have led to smaller, more efficient models without significant sacrifices in accuracy or speed.
-These advancements make them more attractive for modern applications, where both model size and computational efficiency are critical.
-However, despite being the latest addition, YOLOv12 did not follow this trend, ranking third in terms of model size and computational efficiency.
-IV-B4 Overall Performance
-Considering accuracy, speed, size, and GFLOPs, YOLO11m, YOLO11n, YOLO11s, and YOLOv10s emerged as the most consistent performers.
-These models demonstrated a strong balance between high accuracy, low processing time, and efficient disk usage, making them suitable for a wide range of applications where both speed and accuracy are crucial.
-YOLO11 models, in particular, benefited from recent enhancements such as the C3k2 block and C2PSA module, which contributed to their superior performance in these areas.
-The incorporation of these innovations allowed the YOLO11 models to handle larger datasets more effectively without compromising on speed or efficiency.
-On the other hand, YOLOv9e, YOLOv5ux, and YOLOv3u showed poor results across all metrics.
-These models were computationally inefficient, struggling with high GFLOPs counts and slower processing times.
-YOLOv3u, in particular, ranked poorly due to its reliance on the outdated Darknet-53 architecture, which is no longer optimized for modern object detection tasks.
-Despite its larger size, it failed to offer the necessary accuracy, making it less suitable for practical applications.
-Similarly, YOLOv5ux and YOLOv9e showed poor performance in terms of both speed and accuracy, which further limits their real-world usability.
-YOLO11 models were clearly the best performers overall, owing to their modern design and architectural improvements, making them highly suitable for a wide range of use cases.
-YOLOv10 models, although slightly underperforming in accuracy, excelled in computational efficiency due to their use of the One-to-One head for predictions.
-This design allowed YOLOv10 to process data more quickly and efficiently, making it ideal for time-sensitive applications where speed is prioritized over perfect accuracy.
-While YOLOv9 models showed underperformance in terms of computational efficiency, they still remain competitive with YOLOv10 and YOLO11 in terms of accuracy, particularly due to their PGI integration, which helps in enhancing detection precision.
-This makes YOLOv9 a viable option for applications where high accuracy is prioritized over speed.
-Concerning the YOLOv12 family, it showcased mediocre results, underperforming compared to its predecessors, particularly YOLOv9 and YOLO11, in both accuracy and speed.
-Despite introducing innovative components like the Area Attention Module (A2), Residual Efficient Layer Aggregation Networks (R-ELAN), and FlashAttention, these architectural changes contributed to increased computational overhead and slower inference times.
-The Area Attention Module segmented the feature map to maintain a large receptive field but limited contextual understanding, leading to lower accuracy.
-Meanwhile, R-ELAN improved feature aggregation but introduced latency due to its complex residual connections.
-FlashAttention aimed to optimize memory access but added processing overhead, further slowing down the model.
-Although YOLOv12 leveraged convolution operators for efficiency, it failed to offset the complexity introduced by its attention mechanisms.
-Overall, YOLOv12 struggled to balance speed and accuracy, making it less favorable compared to the more efficient and accurate YOLOv9, YOLOv10, and YOLO11 models.
-In addition, YOLOv8 and YOLOv5u achieved competitive results, surpassing YOLOv3u in terms of accuracy.
-This improvement can be attributed to their more recent architectures, which allowed them to handle modern detection tasks better.
-However, despite their higher accuracy compared to YOLOv3u, they still fell short when compared to newer models like YOLOv9, YOLOv10, and YOLO11 in both accuracy and overall performance.
-While YOLOv8 and YOLOv5u had faster processing times than YOLOv9, their overall performance was still inferior to that of the YOLO11 and YOLOv10 models, which maintained a better balance between speed and accuracy.
-IV-B5 The Rise of YOLO11 Over YOLOv8
-Although YOLOv8 [27] has been the algorithm of choice for its versatility in tasks such as pose estimation, instance segmentation, and oriented object detection (OBB), YOLO11 [28] has now emerged as a more efficient and accurate alternative.
-With its ability to handle the same tasks while offering improved contextual understanding and better architectural modules, YOLO11 sets a new standard in performance, surpassing YOLOv8 in both speed and accuracy across various applications.
-This paper contributes by providing empirical evidence of YOLO11’s superior performance in both accuracy and computational efficiency.
-Our findings establish YOLO11 as a clear advancement in the YOLO family, setting a new benchmark for object detection.
-By incorporating novel modules, YOLO11 addresses both the need for high accuracy and low-latency processing, offering valuable insights for deploying YOLO-based models in practical, real-world applications.
-IV-B6 YOLOv9: Efficient Yet Slow
-Although YOLOv9 [70] excels in GFLOPs count and model size, outperforming many other models, it paradoxically ranks as one of the slowest in terms of inference speed.
-This unusual combination of high computational efficiency but low speed can be attributed to YOLOv9’s architectural choices, such as the use of Programmable Gradient Information (PGI) and Generalized Efficient Layer Aggregation Network (GELAN), which optimize computational performance but add significant latency during inference.
-This paper contributes by highlighting the trade-offs in YOLOv9’s design, showing that while its smaller model size and lower GFLOPs count make it ideal for resource-constrained environments, its slow inference times make it unsuitable for time-sensitive applications.
-Our findings suggest that YOLOv9’s performance is best suited for scenarios where computational efficiency is prioritized over speed, offering valuable insights for selecting models based on specific application needs.
-IV-B7 YOLOv12: Ambitious but Underwhelming
-YOLOv12, despite being the latest addition to the YOLO family, delivered subpar results, failing to outperform its predecessors, particularly YOLOv9, YOLOv10, and YOLO11, in both accuracy and speed.
-This underperformance is primarily due to its complex architectural choices, including the Area Attention Module (A2), Residual Efficient Layer Aggregation Networks (R-ELAN), and FlashAttention, which introduced significant computational overhead.
-Although these modules aimed to enhance contextual understanding and memory efficiency, they ultimately compromised speed and accuracy due to increased latency and reduced spatial awareness.
-This work contributes by revealing the trade-offs inherent in YOLOv12’s ambitious design, demonstrating that its added complexity did not translate into superior performance.
-IV-B8 Dataset Size
+YOLOv3u 7 1 4 7
+YOLOv5u 6 4 4 5
+YOLOv8 5 5 7 6
+YOLOv9 3 6 6 4
+YOLOv10 4 2 3 1
+YOLO11 1 3 1 2
+YOLOv12 1 7 2 3
+5.1 Performance Analysis by Model Scale
+Nano Models
+Among nano models, YOLOv9t ranked highest in accuracy and speed, benefiting from its lightweight GELAN architecture, which optimizes parameter efficiency while maintaining strong detection capabilities.
+However, YOLOv10n, despite its computational efficiency, exhibited lower accuracy due to its reliance on the One-to-One head and NMS-Free training, which struggled with densely packed objects.
+Small Models
+YOLOv12s emerged as the best small model in terms of accuracy, likely due to its R-ELAN-based feature aggregation.
+However, it performed poorly in speed, as the added complexity of Area Attention (A2) increased inference time.
+In contrast, YOLOv10s and YOLO11s provided a balanced performance, with YOLO11s leveraging the C3k2 block for improved feature extraction and YOLOv10s excelling in speed due to its efficient head design.
+Medium Models
+YOLO11m achieved the best balance of accuracy and efficiency in this category, owing to C2PSA, which enhances spatial feature capture.
+YOLOv10m outperformed all models in terms of computational efficiency but had lower accuracy due to its reliance on NMS-Free training.
+Meanwhile, YOLOv9m, while achieving strong accuracy, ranked lower in efficiency due to its reliance on gradient-based optimizations, which added computational overhead.
+Large Models
+YOLO11l outperformed other large models in accuracy and speed, thanks to the integration of C2PSA and refined convolutional operations.
+In contrast, YOLOv9c and YOLOv12l struggled with efficiency, with YOLOv9c being computationally expensive and YOLOv12l suffering from latency due to FlashAttention’s memory overhead.
+Extra-Large Models
+YOLO11x provided the best balance of accuracy and computational efficiency, with its C3k2-based optimization reducing inference overhead.
+However, YOLOv9e and YOLOv12x exhibited slower speeds and inefficient GFLOPs utilization, highlighting the limitations of their respective architectures when scaling up.
+5.2 Performance Analysis by Family
+YOLOv12
+YOLOv12 delivered strong accuracy results, ranking first alongside YOLO11 in accuracy across models.
+This performance can be attributed to its integration of the Area Attention Module (A2) and Residual Efficient Layer Aggregation Networks (R-ELAN), which enhanced feature extraction and contextual understanding.
+However, this came at the cost of computational efficiency, as the added complexity introduced significant latency and reduced processing speed.
+FlashAttention, while designed to optimize memory access, further contributed to increased inference time.
+While YOLOv12 excelled in accuracy, its slow processing speed and higher computational demands limited its overall practicality, highlighting the trade-off between architectural complexity and real-world efficiency.
+YOLO11
+The YOLO11 family consistently ranked among the best due to its C3k2 block and C2PSA, enhancing efficiency and contextual understanding by ranking first in terms of accuracy and GFLOPs count and among the first in terms of speed and size.
+These innovations enabled YOLO11 to achieve high accuracy with low computational overhead, making it a well-balanced choice.
+While YOLOv8 was widely used for various tasks such as Pose Estimation and OBB, YOLO11 has emerged as a superior alternative, offering better feature extraction, inference speed, and accuracy.
+This study establishes YOLO11 as a new benchmark, demonstrating its high detection accuracy and low-latency processing, and various capabilities, making it ideal for real-world applications requiring both speed and precision.
+YOLOv10
+YOLOv10 models were highly efficient, particularly in speed and size, due to the One-to-One head and NMS-Free training.
+However, their accuracy was subpar compared to YOLOv9, YOLO11, and YOLOv12, particularly in scenarios requiring precise bounding box refinement, such as the Traffic Signs dataset.
+YOLOv9
+Despite strong accuracy due to PGI and GELAN, YOLOv9 models suffered from slow inference times, limiting their real-time applicability.
+Their architectural choices prioritized precision over efficiency, making them suitable for applications where accuracy is paramount but speed is less critical.
+YOLOv8, YOLOv5u, and YOLOv3u
+While YOLOv8, YOLOv5usurpassed YOLOv3u in accuracy, they lagged behind newer models in overall efficiency.
+YOLOv8 introduced the C2k module for improved feature extraction, but its higher computational cost made it less efficient than more recent architectures.
+With YOLO11 offering superior accuracy, speed, and efficiency, YOLOv8 is no longer the go-to model for real-world applications, as YOLO11 provides a better balance between performance and computational overhead.
+Meanwhile, YOLOv5u improved on YOLOv3u but lacked modern enhancements like self-attention, limiting its scalability.
+Overall, the rankings highlight the dominance of YOLO11 as the most balanced family, with YOLOv10 excelling in efficiency and YOLOv9 maintaining strong accuracy at the cost of speed.
+YOLOv12, despite its ambitious design and exceptional accuracy, failed to outperform its predecessors due to increased latency, reinforcing the importance of balancing architectural complexity with real-world feasibility.
+5.3 Dataset Size
 The size of the dataset significantly influences the performance of YOLO models.
@@
 Conversely, small models like YOLOv9t and YOLOv9s performed best on the Africa Wildlife dataset, showcasing the effectiveness of small-scaled models when handling limited datasets.
-IV-B9 Impact of Training Datasets
-The performance of YOLO models is influenced by the training datasets used, as shown in Tables VI, VII, and VIII.
+5.4 Impact of Training Datasets
+The performance of YOLO models is influenced by the training datasets used, as shown in Figures 12, 16, and 20.
 Different datasets yield varying results and top performers, indicating that dataset complexity affects algorithm performance.
 This underscores the importance of using diverse datasets during benchmarking to obtain comprehensive results on the strengths and limitations of each model.
-This discussion highlights the need for a balanced consideration of accuracy, speed, and model size when selecting YOLO models for specific applications.
-The consistent performance of YOLO11 models across various metrics makes them highly recommended for versatile situations where accuracy and speed are essential.
-Meanwhile, YOLOv10 models can perform similarly while achieving faster processing times and with smaller model sizes.
-Additionally, YOLOv9 can deliver comparable results in terms of accuracy but sacrifices speed, making it suitable for applications where precision is prioritized over rapid processing.
-V Conclusion
+Additionally, this highlights the need for a balanced consideration of accuracy, speed, and model size when selecting YOLO models for specific applications.
+6 Real-Life Applications
+Table 5: YOLO Model Recommendations for Real-Life Applications
+Scenario Recommended YOLO Models Application Examples
+Computationally Constrained Environments YOLOv10(n/s/m) Drone surveillance, embedded systems, battery-powered devices [29, 80, 25, 27, 9, 8]
+YOLO11(n/s/m)
+Real-Time Monitoring and Rapid Response YOLOv10(n/s/m) Autonomous vehicles, disaster response, traffic monitoring [29, 80, 25, 75, 27, 9, 8]
+YOLO11(n/s/m)
+Detection of Small Objects YOLO11(m/l/x) Wildlife tracking, satellite imaging of ships and small vehicles [75, 27, 9, 8]
+YOLOv12(m/l/x)
+Detection of Large Objects YOLOv8(s/m/l/x) Satellite imaging (buildings, forests, urban development) [83, 55, 43]
+YOLOv9(t/s/m/c/e)
+Overlapping and Densely
+Packed Objects
+YOLOv9(t/s/m/c/e)
+Traffic intersections,
+wildlife in dense foliage
+Oriented and Rotated Object Detection YOLO11 (OBB Support) Satellite imagery, aerial monitoring of ships [46, 21]
+YOLOv12 (OBB Support)
+Handling Large Datasets YOLO11(m/l/x) City-wide monitoring, biodiversity mapping [5, 76]
+YOLOv12(m/l/x)
+Handling Small Datasets YOLOv9(t/s/m/c)
+Anti-poaching monitoring,
+niche industrial tasks [1, 64, 84]
+The advancements in YOLO models have enabled their application across diverse real-world scenarios, each with distinct demands for model performance, such as computational efficiency, accuracy, and adaptability to varying object complexities.
+Based on the benchmark of various YOLO models and related papers, this section highlights how different models excel under specific scenarios, making them suitable for deployment in various tasks across industries, as shown in Table 5.
+6.1 Computationally Constrained Environments
+In scenarios with limited computational resources, such as drones, embedded systems, and battery-powered devices, models must be lightweight and efficient to avoid draining power or overloading hardware.
+The nano, small, and medium-sized models of the YOLO11 and YOLOv10 families are optimal choices due to their high accuracy, compact model sizes, low inference times, and low GFLOPs count leading to minimal resource consumption, as also seen in related papers [29, 80, 25, 27, 9, 8].
+6.2 Real-Time Monitoring and Rapid Response
+For applications requiring real-time detection, such as autonomous driving, surveillance, and disaster response, models need to balance accuracy and speed.
+As illustrated by this study and related research papers [29, 80, 25, 75, 27, 9, 8], the nano, small, and medium-sized YOLO11 and YOLOv10 models demonstrate strong performance in maintaining fast inference while accurately detecting objects across various scales.
+6.3 Detection of Small and Large Objects
+Object detection models often face the challenge of accurately identifying small and large objects.
+Small objects, such as wildlife or distant vehicles, require models capable of capturing fine spatial details.
+Medium, large, and extra-large-sized YOLO11 and YOLOv12 models excel in these scenarios, as illustrated in our paper and related work [75, 27, 9, 8].
+Meanwhile, all YOLOv9 and YOLOv8 models offer robust performance for detecting large objects, making them ideal for scenarios such as satellite imaging of buildings, forests, and urban development [83, 55, 43].
+6.4 Overlapping and Densely Packed Objects
+Detecting objects that overlap or appear densely packed is a major challenge in traffic intersections or wildlife monitoring scenarios.
+In these cases, models must efficiently distinguish between objects and avoid false positives.
+All YOLOv9 models are highly effective in handling overlapping objects, as seen in this benchmark and related work [46, 21].
+This suits them, particularly for applications like detecting animals in dense foliage or distinguishing closely packed traffic signs.
+6.5 Oriented Object Detection
+Objects often appear at varying angles or orientations in satellite imaging and certain surveillance applications, requiring models with specialized detection capabilities.
+The YOLO11 and YOLOv12 families’ support for OBB enables effective detection of rotated objects, achieving higher accuracy than YOLOv8 models.
+6.6 Handling Large and Very Small Datasets
+Large-scale datasets with diverse object categories require models that generalize well without overfitting.
+Medium, large, and extra-large YOLO11 and YOLOv12 models excel in such scenarios, as showcased in our paper and related studies [5, 76].
+For tiny datasets, where overfitting is a concern, all YOLOv9 models are optimal choices.
+Their compact sizes and efficient parameter usage enable effective learning from limited data [1, 64, 84].
+7 Conclusion
 This benchmark study thoroughly evaluates the performance of various YOLO algorithms.
-It pioneers a comprehensive comparison of YOLO11 against its predecessors, evaluating their performance across three diverse datasets: Traffic Signs, African Wildlife, and Ships and Vessels.
-The datasets were carefully selected to encompass a wide range of object properties, including varying object sizes, aspect ratios, and object densities.
+It pioneers a comprehensive comparison of YOLOv12 against its predecessors and assesses their performance across three diverse datasets: Traffic Signs, African Wildlife, and Ships and Vessels.
+The datasets were carefully selected to encompass various object properties, including varying object sizes, aspect ratios, and object densities.
 We showcase the strengths and weaknesses of each YOLO version and family by examining a wide range of metrics such as Precision, Recall, Mean Average Precision (mAP), Processing Time, GFLOPs count, and Model Size.
 Our study addresses the following key research questions:
+  • 1.
     Which YOLO algorithm demonstrates superior performance across a comprehensive set of metrics?
+  • 2.
     How do different YOLO versions perform on datasets with diverse object characteristics?
-  • What are the specific strengths and limitations of each YOLO version, and how can these insights inform the selection of suitable algorithms for various applications?
+  • 3.
+    What are each YOLO version’s specific strengths and limitations, and how can these insights inform the selection of suitable algorithms for various applications?
+  • 4.
+    Which YOLO models are best suited for specific real-world scenarios?
 In particular, the YOLO11 family emerged as the most consistent, with YOLO11m striking an optimal balance between accuracy, efficiency, and model size.
@@
 These findings provide valuable insights for industry and academia, guiding the selection of the most suitable YOLO algorithms and informing future developments and enhancements.
-YOLOv12, despite being the newest addition to the YOLO family, delivered underwhelming performance.
+Despite being the newest addition to the YOLO family, YOLOv12 delivered an underwhelming performance.
 Its ambitious architectural changes, including the Area Attention Module and R-ELAN, introduced complexity without translating into clear advantages.
-The model struggled to balance accuracy and speed, highlighting the challenges of integrating advanced attention mechanisms effectively within the YOLO framework.
+The model struggled to balance accuracy and speed, highlighting the challenges of effectively integrating advanced attention mechanisms within the YOLO framework.
 While the evaluated algorithms show promising performance, there is room for improvement.
@@
 Our future work will address these gaps and explore enhancements to maximize overall efficiency and real-world applicability.
+Declarations
+  • 1.
+    Funding: Not applicable
+  • 2.
+    Conflict of interest/Competing interests: The authors declare no conflict of interest.
+  • 3.
+    Ethics approval and consent to participate: Not applicable
+  • 4.
+    Consent for publication: The authors confirm that all necessary permissions have been obtained for the publication of this work, including the use of copyrighted material and any personal information or sensitive data.
+  • 5.
+    Data availability: Data is available on GitHub and will be publicly available upon publication of the final version
+  • 6.
+    Materials availability: Not applicable
+  • 7.
+    Code availability: Source code is available on GitHub and will be publicly available upon publication of the final version.
+  • 8.
+    Author contribution: All authors contributed to the study conception and design.
+Material preparation, data collection, and analysis were performed by Nidhal Jegham and Chan Young Koh.
+The first draft of the manuscript was written by Nidhal Jegham, and all authors commented on and edited previous versions of the manuscript.
+All authors read and approved the final manuscript.
+Dr.
+Abdeltawab Hendawi supervised the work of this study.